### v4

# Equiangular lines via matrix projection

Abstract
In 1973, Lemmens and Seidel posed the problem of determining the maximum number of equiangular lines in $\mathbb{R}^{r}$ with angle $\arccos(\alpha)$ and gave a partial answer in the regime $r\leq 1/\alpha^{2}-2$. At the other extreme where $r$ is at least exponential in $1/\alpha^{2}$, recent breakthroughs have led to an almost complete resolution of this problem. In this paper, we introduce a new method for obtaining upper bounds which unifies and improves upon previous approaches, thereby yielding bounds which bridge the gap between the aforementioned regimes and are best possible either exactly or up to a small multiplicative constant. Our approach relies on orthogonal projection of matrices with respect to the Frobenius inner product and as a byproduct, it yields the first extension of the Alon–Boppana theorem to dense graphs, with equality for strongly regular graphs corresponding to families of $\binom{r+1}{2}$ equiangular lines in $\mathbb{R}^{r}$. Applications of our method in the complex setting will be discussed as well.
1 Introduction
Given $n$ lines $l_{1},\ldots,l_{n}$ passing through the origin in $\mathbb{R}^{r}$, we say that they are equiangular if there exists a common angle $\theta\in(0,\pi/2]$ such that for all $i\neq j$, the acute angle between $l_{i}$ and $l_{j}$ is $\theta$. If we choose a unit vector $v_{i}$ along each line $l_{i}$, an equivalent definition is that $|\left\langle v_{i},v_{j}\right\rangle|=\cos(\theta)$ for all $i\neq j$, and this latter definition extends to complex lines in $\mathbb{C}^{r}$. Large sets of equiangular lines have been studied for over 75 years, arising naturally in a wide variety of different areas including elliptic geometry [50, 13], the theory of polytopes [24], frame theory [53, 20, 51], the theory of Banach spaces [10, 62, 37, 30], and perhaps more surprisingly, quantum information theory [91, 77, 43, 42, 21, 41] with connections to algebraic number theory and Hilbert’s twelfth problem [11, 5, 6, 4]. Moreover, determining the maximum number of equiangular lines in $\mathbb{R}^{r}$ is considered to be one of the founding problems of algebraic graph theory [45] and it is well known due to the absolute bound (see [67]) that the answer cannot be larger than $\binom{r+1}{2}$, which is tight up to a multiplicative constant.
In this paper, we will focus on the more refined question of determining $N^{\mathbb{R}}_{\alpha}(r)$, the maximum number of equiangular lines in $\mathbb{R}^{r}$ with common angle $\arccos(\alpha)$. This problem was first posed by Lemmens and Seidel [67] in 1973, who gave a good partial answer when $r\leq 1/\alpha^{2}-2$ by proving the relative bound, for which many tight constructions are known [35]. On the other hand, prior to this work, our understanding of the complementary regime $r\geq 1/\alpha^{2}-2$ has been much more limited. Recently, several breakthroughs have led to an almost complete solution of this problem when $r$ is very large relative to $1/\alpha$. The first result in this direction was due to Bukh [14], which we subsequently improved in a joint work with Dräxler, Keevash, and Sudakov [7], in particular showing that $N^{\mathbb{R}}_{\alpha}(r)=\Theta(r)$ when $r$ is at least exponential in $1/\alpha^{2}$. Jiang and Polyanskii [59] obtained further refinements, culminating in the work of Jiang, Tidor, Yao, Zhang, and Zhao [61] who determined $N^{\mathbb{R}}_{\alpha}(r)$ completely when $r$ is at least doubly exponential in $k/\alpha$, where $k=k\left(\frac{1-\alpha}{2\alpha}\right)$ is the corresponding spectral radius order (see the end of Section 1.4 for a definition). However, it is not possible to extend these recent results to the remaining regime $1/\alpha^{2}-2\leq r\leq 2^{\Theta(1/\alpha^{2})}$ because they all rely on the framework we introduced in [7], which uses Ramsey’s theorem from graph theory. Note that the only other bounds known for general $\alpha$ in this regime were also obtained more recently by Yu [90] and Glazyrin and Yu [44] using semidefinite programming and properties of Gegenbauer polynomials.
In this paper, we overcome the bottleneck of the Ramsey-theoretic approach by introducing a new method via orthogonal projection of matrices with respect to the Frobenius inner product, thereby allowing us to significantly extend or improve almost all of the aforementioned results. Indeed, we obtain upper bounds on $N^{\mathbb{R}}_{\alpha}(r)$ for all $r\geq 1/\alpha^{2}-2$, which are best possible [^1] When we say best possible, we mean either for all $\alpha$ or for infinitely many $\alpha\rightarrow 0$, depending on the regime. either exactly or up to a small multiplicative constant. As a consequence, we determine that $N^{\mathbb{R}}_{\alpha}(r)=\Theta(r)$ when $r\geq\Omega\left(1/\alpha^{4}\right)$ and for infinitely many $\alpha\rightarrow 0$, we determine that $N^{\mathbb{R}}_{\alpha}(r)=\Theta\left(1/\alpha^{4}\right)$ when $C/\alpha^{2}\leq r\leq O\left(1/\alpha^{4}\right)$ for some constant $C$. We also extend a bound obtained implicitly by Jiang, Tidor, Yao, Zhang, and Zhao [61] from $r\geq 2^{2^{\Theta\left(1/\alpha\right)}}$ to $r\geq 2^{1/\alpha^{\Theta(1)}}$ and thereby determine $N^{\mathbb{R}}_{\alpha}(r)$ completely when $r$ is at least doubly exponential in $k\log\left(1/\alpha\right)$, where $k$ is the spectral radius order [^2] Note that $k$ is always at least $\frac{1-3\alpha}{2\alpha}$, so we only obtain a modest extension of the main result of [61].. Moreover, we also observe that when $r\leq 1/\alpha^{2}-2$, our projection method yields a simple proof of the relative bound and more generally, it can be used to obtain a refinement of the first Welch bound [89] from coding theory. It follows that our approach gives a unified framework for deriving all of the best upper bounds on $N^{\mathbb{R}}_{\alpha}(r)$ for general $\alpha$.
Since the unit vectors which span a family of real equiangular lines have pairwise inner product $\alpha$ or $-\alpha$, this information can be encoded in a corresponding graph. Conversely, given any $d$-regular graph whose adjacency matrix has second largest eigenvalue $\lambda_{2}$ satisfying $d-\lambda_{2}\leq(1-\varepsilon)n/2$ where $\varepsilon>0$, there exists a family of equiangular lines which corresponds to it. If we assume for the sake of simplicity that we are in this setting and $\varepsilon$ is a constant, then roughly speaking, our main results for equiangular lines rely on the novel lower bound
$\lambda_{2}\geq\Omega\left(d^{1/3}\right).$
We derive this bound, as well as the related bound $\lambda_{2}\geq\Omega\left(|\lambda_{n}|^{1/2}\right)$ (see Theorem 1.15 for precise statements), by orthogonally projecting specially chosen matrices onto the span of the linear projection matrices $v_{1}v_{1}^{\intercal},\ldots,v_{n}v_{n}^{\intercal}$ corresponding to the given lines. Furthermore, these bounds are tight for strongly regular graphs corresponding to families of equiangular lines meeting the absolute bound in $\mathbb{R}^{r}$. Prior to this work, such bounds were only known for sufficiently sparse graphs (diameter $\geq 4$) via the usual Alon–Boppana theorem and so, our results may be seen as the first extension of this celebrated theorem to dense graphs. Furthermore, our results also rely on a bootstrapping argument using new variants of the usual Alon–Boppana theorem (see Lemma 3.13 and the proof of Lemma 3.14).
Remark 1.1.
After an earlier version of this paper was put on arXiv, Ihringer [57] obtained another proof of our inequality $\lambda_{2}\geq\Omega\left(|\lambda_{n}|^{1/2}\right)$, interpreting it as a variant of the Krein bound for regular graphs. Furthermore, Räty, Sudakov, and Tomon [76] (see also [8] for a short exposition) very recently gave alternative proofs of our new variants of the Alon–Boppana theorem for regular graphs. In particular, they noted that the bound $\lambda_{2}\geq\Omega\left(d^{1/3}\right)$ follows from $\lambda_{2}\geq\Omega\left(|\lambda_{n}|^{1/2}\right)$ and moreover, they pointed out the existence of certain infinite families of strongly regular graphs which imply that both bounds are asymptotically tight when $d=(1-o(1))(2n)^{3/4}$ or tight up to a multiplicative constant when $d=\Theta(n)$.
Note that when trying to obtain bounds on real equiangular lines, the corresponding graph need not be regular and so the interested reader is advised to look at Section 3 in order to see how the aforementioned results are appropriately generalized and implemented. Indeed, our approach actually relies on applying the projection method together with a switching argument in order to obtain control over the largest eigenvalue of a corresponding Gram matrix and to obtain precise bounds on the entries of the corresponding eigenvector. If the largest eigenvalue is at most $\frac{1-\alpha^{2}}{2\alpha^{2}}$, then it is easy to see that the bound of Yu [90] follows and otherwise, we obtain a bound of $O\left(1/\alpha^{3}\right)$ on the maximum degree of a corresponding graph. We then apply new variants of the usual Alon–Boppana theorem in order to bootstrap this bound to $O\left(1/\alpha^{2}\right)$. This maximum degree bound is then combined with combinatorial/spectral arguments as in [7, 61].
We also consider the more general question of determining $N^{\mathbb{C}}_{\alpha}(r)$, the maximum number of complex equiangular lines in $\mathbb{C}^{r}$ with common Hermitian angle $\arccos(\alpha)$ (see Section 1.2). Similarly to the real case, there is an absolute bound of $r^{2}$ and the relative bound holds for $r\leq 1/\alpha^{2}-1$, but no other bounds are known in the complementary regime $r\geq 1/\alpha^{2}-1$. Our approach generalizes to this setting, allowing us to obtain the first such bounds. In particular, if $r\geq\left(1-\Theta(\alpha)\right)/\alpha^{3}$ then we show that $N^{\mathbb{C}}_{\alpha}(r)\leq O(r/\alpha)$. Otherwise, if $r\leq\left(1-\Theta(\alpha)\right)/\alpha^{3}$ then by generalizing the bound of Yu [90] to complex lines, we conclude that $N^{\mathbb{C}}_{\alpha}(r)\leq(1/\alpha^{2}-1)^{2}$ for all $1/\alpha^{2}-1\leq r\leq(1-\Theta(\alpha))/\alpha^{3}$, with equality if and only if the given complex lines lie in a subspace of dimension $d=1/\alpha^{2}-1$ and they meet the absolute bound of $d^{2}$. Note that an important conjecture in quantum information theory due to Zauner (1.10) states that such families of $d^{2}$ lines in $\mathbb{C}^{d}$ exist for all $d\in\mathbb{N}$ and thus, our result provides a potential alternative route to establishing this conjecture.
Our projection method relies on the simple fact that in an inner product space, the length of a vector cannot increase after an orthogonal projection onto a subspace. This argument has the benefit of yielding tight bounds for collections of lines meeting a corresponding absolute bound, since in such cases the projection becomes the identity map. For this paper, we will take our inner product space to be a space of matrices, i.e. 2-tensors, but we note that by considering higher order symmetric tensors with respect to the Frobenius inner product, one can derive the Gegenbauer polynomials and obtain the classical linear programming bounds for spherical codes due to Delsarte, Goethals, and Seidel [28, 29] (see the discussion at the end of Section 2). Nonetheless, we do not need the full power of such an approach for the results in this paper. Indeed, our projection arguments can also be derived via a special case of the Schur product theorem, i.e. the fact that the entry-wise square of a positive semidefinite matrix is still positive semidefinite (see Remark 2.3).
Structure of the paper
This paper is organized as follows. In Section 1.1, Section 1.2, and Section 1.3 we give more detailed introductions to and summaries of our main results on real equiangular lines, complex equiangular lines, and eigenvalues of regular graphs, respectively. Section 1.4 contains relevant notation and definitions that are used throughout the paper. In Section 2, we discuss how orthogonal projection of matrices can be used to derive the geometric inequalities which underly our results, how this approach relates to the Delsarte linear programming method, and as an additional application, obtain an improvement to the first Welch bound. Then in Section 3, Section 4, and Section 5 we derive our main results for real equiangular lines, complex equiangular lines, and regular graphs, respectively. Finally, in Section 6 we give some concluding remarks and directions for future research.
1.1 Real equiangular lines
Let $N^{\mathbb{R}}(r)$ be the maximum number of equiangular lines in $\mathbb{R}^{r}$. It is not hard to show that $N^{\mathbb{R}}(2)=3$, with the optimal configuration being familiar to anyone who has cut a pizza pie into 6 equal slices using 3 cuts. In 1948, Haantjes [49] showed that $N^{\mathbb{R}}(3)=N^{\mathbb{R}}(4)=6$, with an optimal configuration coming from the 6 diagonals of a regular icosahedron. The question of determining $N^{\mathbb{R}}(r)$ for an arbitrary $r$ was first formally posed in 1966 by van Lint and Seidel [69], and a few years later Gerzon (see [67]) proved the absolute bound
$N^{\mathbb{R}}(r)\leq\binom{r+1}{2}.$
This bound is known to be tight for $r=2,3,7,23$, with the constructions for $r=7$ and $r=23$ being based on the $E_{8}$ lattice in $\mathbb{R}^{8}$ and the Leech lattice in $\mathbb{R}^{24}$, see [67] and references therein. Surprisingly, we do not know if there are any other $r$ for which this is the case. Moreover, the order of magnitude of $N^{\mathbb{R}}(r)$ was not even known until 2000, when de Caen [17] gave a construction showing that $N^{\mathbb{R}}(r)\geq\frac{2}{9}(r+1)^{2}$ for infinitely many $r$.
In 1973, Lemmens and Seidel [67] proposed to study the more refined quantity $N^{\mathbb{R}}_{\alpha}(r)$, the maximum number of equiangular lines in $\mathbb{R}^{r}$ with common angle $\arccos(\alpha)$, where $\alpha\in[0,1)$. They proved the relative bound
$N^{\mathbb{R}}_{\alpha}(r)\leq\frac{1-\alpha^{2}}{1-\alpha^{2}r}r\text{ for all }r<1/\alpha^{2},$
which is known to be tight if and only if the unit vectors $v_{1},\ldots,v_{n}$ which span the lines form a tight frame, i.e. $\sum_{i=1}^{n}{v_{i}v_{i}^{\intercal}=\frac{n}{r}I}$. Equiangular tight frames correspond to strongly regular graphs with certain parameters and subject to some divisibility conditions, they are known to exist for many different values of $r$ and $n$, see the survey of Fickus and Mixon [35]. Lemmens and Seidel [67] also showed that a construction meeting the absolute bound must have $r=1/\alpha^{2}-2$, so that the relative bound can be seen as a refinement of the absolute bound when $r\leq 1/\alpha^{2}-2$. They were particularly interested in the case where $1/\alpha$ is an odd integer due to a result of Neumann (see [67]), who showed that if this is not the case, then $N^{\mathbb{R}}_{\alpha}(r)\leq 2r$.
More recently, Bukh [14] showed that $N^{\mathbb{R}}_{\alpha}(r)$ is linear in $r$ when $\alpha$ is fixed, as well as conjecturing the asymptotic value of $N^{\mathbb{R}}_{\alpha}(r)$ as $r\rightarrow\infty$ whenever $1/\alpha$ is an odd integer. Together with Dräxler, Keevash, and Sudakov [7], we significantly improved this result, showing that there exists a positive constant $C$ such that $N^{\mathbb{R}}_{\alpha}(r)\leq 2r-2$ for all $r\geq 2^{C/\alpha^{2}}$, with equality if and only if $\alpha=1/3$. Some of the ideas we proposed were further clarified and extended by Jiang and Polyanskii [59], who also generalized Bukh’s conjecture to any $\alpha$. Using the methods developed in [7], together with a new bound on the maximum multiplicity of the second eigenvalue of a connected graph, Jiang, Tidor, Yao, Zhang, and Zhao [61] were able to verify this conjecture in a strong form. To state their result, we first define the spectral radius order $k(\lambda)$ to be the least number of vertices in a graph having spectral radius $\lambda$ (we set $k(\lambda)=\infty$ if there is no such graph). If the spectral radius order $k=k\left(\frac{1-\alpha}{2\alpha}\right)<\infty$, it was shown in [61] that there exists a positive constant $C$ such that $N^{\mathbb{R}}_{\alpha}(r)=\left\lfloor\frac{k}{k-1}(r-1)\right\rfloor\text{ for all }r\geq 2^{2^{Ck/\alpha}}$.
All of these recent results apply Ramsey’s theorem in order to obtain a bound on the maximum degree of a corresponding graph which is exponential in $1/\alpha$. Therefore, such methods stop working when $r$ is much smaller than $2^{1/\alpha}$, in which case the only known bounds were obtained using semidefinite programming and Gegenbauer polynomials. Yu [90] showed that $N^{\mathbb{R}}_{\alpha}(r)\leq\binom{1/\alpha^{2}-1}{2}$ for $161/\alpha^{2}-2\leq r\leq 3/\alpha^{2}-$ 16 and $\alpha\leq 1/3$, which can be seen as a refinement of the absolute bound when $r\geq 1/\alpha^{2}-2$ and later Glazyrin and Yu [44] characterized the case of equality, as well as proving the universal bound $N^{\mathbb{R}}_{\alpha}(r)\leq\left(\frac{2}{3\alpha^{2}}+\frac{4}{7}\right)r+2$ for all $\alpha\leq 1/3$. Note that Yu’s bound is best possible when the absolute bound is met in some dimension, but it only applies when $r$ is very close to $1/\alpha^{2}$, and in view of our linear bound $22r-$ 2 obtained in [7], Glazyrin and Yu’s universal bound is off by a factor of $1/\alpha^{2}$ when $r$ is large relative to $1/\alpha$.
In this paper, we introduce a new method based on orthogonal projection of symmetric matrices with respect to the Frobenius inner product in order to unify and improve upon previous approaches to obtaining upper bounds on $N^{\mathbb{R}}_{\alpha}(r)$. As a result, we are able to extend our arguments from [7] to the entire range $r\geq 1/\alpha^{2}-2$, thereby obtaining the following theorem which roughly shows that the correct answer is the maximum of our linear bound from [7] and the $\binom{1/\alpha^{2}-1}{2}$ bound of Yu [90]. Moreover, for the regime not covered by the relative bound, i.e. for all $r\geq 1/\alpha^{2}-2$, this theorem significantly extends the bound of Yu and is best possible either exactly or up to a small multiplicative constant.
Theorem 1.2.
Let $10<\alpha<$ 1 and define $r_{\alpha}=\binom{1/\alpha^{2}-1}{2}/(2+3\alpha)$. Then for all $r\leq r_{\alpha}$, we have
$N^{\mathbb{R}}_{\alpha}(r)\leq\binom{1/\alpha^{2}-1}{2}$
with equality only if the corresponding lines span a $(1/\alpha^{2}-2)$-dimensional subspace. Furthermore, for all $r>r_{\alpha}$, we have
$N^{\mathbb{R}}_{\alpha}(r)\leq\left(2+3\alpha\right)r.$
Remark 1.3.
As previously mentioned, it is known that if the absolute bound is met in $\mathbb{R}^{r^{\prime}}$, then the corresponding $\alpha$ must satisfy $r^{\prime}=1/\alpha^{2}-2$, in which case Theorem 1.2 would imply that $N^{\mathbb{R}}_{\alpha}(r)=\binom{1/\alpha^{2}-1}{2}$ for all $\frac{1}{\alpha^{2}}-2\leq r<r_{\alpha}=\frac{1-O(\alpha)}{4\alpha^{4}}$. Conversely, if $N^{\mathbb{R}}_{\alpha}(r)=\binom{1/\alpha^{2}-1}{2}$ for some $\frac{1}{\alpha^{2}}-2\leq r<r_{\alpha}=\frac{1-O(\alpha)}{4\alpha^{4}}$, then Theorem 1.2 implies that $1/\alpha^{2}$ is an integer and the corresponding collection of lines meets the absolute bound in a $(1/\alpha^{2}-2)$-dimensional subspace.
The proof of Theorem 1.2 follows via bounds on the largest eigenvalue of a corresponding Gram matrix, see the beginning of Section 3 for an outline. Moreover, in view of Remark 1.3, improving this theorem for general $\alpha$ when $\frac{1}{\alpha^{2}}-2\leq r\leq\frac{1}{4\alpha^{4}}-\Theta\left(\frac{1}{\alpha^{2}}\right)$ would require showing that the absolute bound can be improved for all sufficiently large dimensions, which is a difficult open problem [^3] The question of whether the absolute bound is tight in a given dimension is equivalent to the existence of a strongly regular graph with certain parameters, see Corollary 5.6 in [88].. On the other hand, when $r$ is much larger than $r_{\alpha}$, we obtain the following improved bounds, in the same way Jiang and Polyanskii [59] refined our approach in [7].
Theorem 1.4.
Let $s\geq 2$ be an integer and let $10<\alpha<$ 1. If $r\gg 1/\alpha^{2s+1}$, then
$N^{\mathbb{R}}_{\alpha}(r)\leq\left(1+\left(\frac{2\alpha}{1-\alpha}\right)^{2}+\frac{1+o(1)}{4\cos^{2}\left(\frac{\pi}{s+2}\right)}\right)r.$
In particular, if $r\geq 1/\alpha^{\omega(1)}$, then $N^{\mathbb{R}}_{\alpha}(r)\leq\left(\frac{5}{4}+\left(\frac{2\alpha}{1-\alpha}\right)^{2}+o(1)\right)r$.
Finally, when $r$ is exponential in $1/\alpha$, we obtain the following extension of the results of Jiang, Tidor, Yao, Zhang, and Zhao [61].
Theorem 1.5.
There exists a constant $C>0$, such that for all $10<\alpha<$ 1 with corresponding spectral radius order $k=k\left(\frac{1-\alpha}{2\alpha}\right)$, we have
  1. 1.
    $N^{\mathbb{R}}_{\alpha}(r)=\left\lfloor\frac{k}{k-1}(r-1)\right\rfloor$ if $k<\infty$ and $r\geq 2^{1/\alpha^{C(k-1)}}$,
  2. 2.
    $N^{\mathbb{R}}_{\alpha}(r)\leq r+\frac{Cr\log(1/\alpha)}{\log{\log{r}}}$ if $k<\infty$ and $2^{1/\alpha^{C}}\leq r<2^{1/\alpha^{C(k-1)}}$,
  3. 3.
    $N^{\mathbb{R}}_{\alpha}(r)\leq r+\frac{Cr\log(1/\alpha)}{\log{\log{r}}}$ if $k=\infty$ and $r\geq 2^{1/\alpha^{C}}$.
Note that for any $\lambda$, the spectral radius order $k(\lambda)\geq\lambda-1$. Moreover, in the special case that $\alpha=\frac{1}{2t-1}$ for some integer $t\geq 2$, we have $k\left(\frac{1-\alpha}{2\alpha}\right)=t$ and thus we obtain the following corollary.
Corollary 1.6.
There exists a constant $C>0$ such that for all $r,k\in\mathbb{N}$ with $k\geq 2$ and $r\geq 2^{k^{Ck}}$,
$N^{\mathbb{R}}_{\frac{1}{2k-1}}(r)=\left\lfloor\frac{k}{k-1}(r-1)\right\rfloor.$
We now summarize the consequences of our new upper bounds Theorem 1.2, Theorem 1.4, and Theorem 1.5, together with the relative bound. In the following, $k=k\left(\frac{1-\alpha}{2\alpha}\right)$ is the corresponding spectral radius, $C$ is a positive constant, and all of the asymptotic bounds are assuming $r\rightarrow\infty$ and $\alpha\rightarrow 0$.
$N^{\mathbb{R}}_{\alpha}(r)\leq\begin{cases}\frac{1-\alpha^{2}}{1-\alpha^{2}r}r&\text{if }r\leq\frac{1}{\alpha^{2}}-2\\[4.0pt] \binom{1/\alpha^{2}-1}{2}&\text{if }\frac{1}{\alpha^{2}}-2<r\leq\frac{1-o(1)}{4\alpha^{4}}\\[4.0pt] \left(2+o(1)\right)r&\text{if }\frac{1-o(1)}{4\alpha^{4}}<r\leq O\left(\frac{1}{\alpha^{5}}\right)\\[4.0pt] \left(1+\frac{1+o(1)}{4\cos^{2}\left(\frac{\pi}{q+2}\right)}\right)r&\text{if }\frac{1}{\alpha^{2q+1}}\ll r\leq O\left(\frac{1}{\alpha^{2q+3}}\right)\text{ for integer }q\geq 2\\[4.0pt] \left(\frac{5}{4}+o(1)\right)r&\text{if }1/\alpha^{\omega(1)}\leq r<2^{1/\alpha^{4C}}\\[4.0pt] \left(1+\frac{C\log(1/\alpha)}{\log{\log{r}}}\right)r&\text{if }2^{1/\alpha^{4C}}\leq r<2^{1/\alpha^{C(k-1)}}\\[4.0pt] \left\lfloor\frac{k}{k-1}(r-1)\right\rfloor&\text{if }2^{1/\alpha^{C(k-1)}}\leq r.\end{cases}$
We also note that Lemmens and Seidel [67] already determined that $N^{\mathbb{R}}_{1/3}(r)=\max\left(28,2r-2\right)$ for all $r\geq 7$ and moreover, they also conjectured that $N^{\mathbb{R}}_{1/5}(r)=\max\left(276,\lfloor 3(r-1)/2\rfloor\right)$ for all $r\geq 23$, which was recently proven by Cao, Koolen, Lin, and Yu [19], building off of the work of Neumaier [72]. In view of these results, we propose the following more general conjecture.
Conjecture 1.7.
For all $10<\alpha<$ 1 with corresponding spectral radius order $k=k\left(\frac{1-\alpha}{2\alpha}\right)<\infty$,
$N^{\mathbb{R}}_{\alpha}(r)\leq\max\left(\binom{1/\alpha^{2}-1}{2},\left\lfloor\frac{k}{k-1}(r-1)\right\rfloor\right).$
Even if this ambitious conjecture is false in general, it would already be very interesting to see if it holds in the case where $1/\alpha$ is an odd integer. Note that Theorem 1.2 establishes a weak version of 1.7 and moreover, if one could significantly improve Jiang, Tidor, Yao, Zhang, and Zhao’s bound [61] on the multiplicity of the second eigenvalue of a connected graph, then 1.7 would follow from our arguments.
We conclude this subsection with some lower bounds for $N^{\mathbb{R}}_{\alpha}(r)$, thereby showing that there exists a constant $C$ such that Theorem 1.2 is best possible up to a small multiplicative constant when $r\geq C/\alpha^{2}$. For all $\alpha\in[0,1)$ and $r\in\mathbb{N}$, it is not hard to see that $N^{\mathbb{R}}_{\alpha}(r)\geq r$. Indeed, if we rotate each standard basis in $\mathbb{R}^{r}$ by the same angle towards the all ones vector $\mathbbm{1}$, we can obtain a collection [^4] More precisely, for $i\in[r]$ let $v_{i}=\left(\sqrt{\frac{1-\alpha}{r}+\alpha}-\sqrt{\frac{1-\alpha}{r}}\right)\mathbbm{1}+\left(\sqrt{1-\alpha}\right)e_{i}$ in $\mathbb{R}^{r}$. Then $\{v_{1},\ldots,v_{r}\}$ is a collection of unit vectors in $\mathbb{R}^{r}$ with pairwise inner product $\alpha$. of $r$ unit vectors with all pairwise inner products $\alpha$. Using this fact, we obtain the following immediate corollary of Theorem 1.2.
Corollary 1.8.
For all $10<\alpha<$ 1 and $r\geq\Omega(1/\alpha^{4})$, we have $N^{\mathbb{R}}_{\alpha}(r)=\Theta(r)$.
In view of the aforementioned result of Neumann, we can only hope to obtain lower bounds which are superlinear in $r$ when $1/\alpha$ is an odd integer. Moreover, it is known that all constructions of $\Omega(r^{2})$ lines in $\mathbb{R}^{r}$ satisfy $r=\Theta(1/\alpha^{2})$. In particular, de Caen’s construction [17] works for all dimensions $r_{t}=3\cdot 2^{2t-1}-1$ where $t\in\mathbb{N}$. Using the fact that $N^{\mathbb{R}}_{\alpha}(r)$ is a non-decreasing function of $r$, we immediately obtain the following corollary of de Caen’s construction and Theorem 1.2.
Corollary 1.9.
There exists a constant $C>0$ and an infinite sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ with $\alpha_{t}\rightarrow 0$ such that for all $\alpha=\alpha_{t}$ and $\frac{C}{\alpha^{2}}\leq r\leq O\left(\frac{1}{\alpha^{4}}\right)$, we have $N^{\mathbb{R}}_{\alpha}(r)=\Theta(1/\alpha^{4})$.
Using Theorem 1.2, together with a family of strongly regular graphs constructed by Metz (see Remark 1.18), we also note that there exists a constant $C$ such that for any $\alpha=\frac{1}{2q-1}$ where $q$ is a prime power, $N^{\mathbb{R}}_{\alpha}(r)=\Theta\left(1/\alpha^{4}\right)$ when $C/\alpha^{3}\leq r\leq O\left(1/\alpha^{4}\right)$.
1.2 Complex equiangular lines
For a pair of 1-dimensional subspaces $U,V\subseteq\mathbb{C}^{r}$, i.e. complex lines through the origin, the quantity $\arccos{|\left\langle u,v\right\rangle|}$ is the same for any choice of unit vectors $u\in U,v\in V$, and so we may define this quantity to be the Hermitian angle between $U$ and $V$, see e.g. [79]. Given $n$ complex lines $l_{1},\ldots,l_{n}$ passing through the origin in $\mathbb{C}^{r}$, we say they are equiangular if there exists $\theta\in(0,\pi/2]$ such that the Hermitian angle between $l_{i}$ and $l_{j}$ is $\theta$ for all $i\neq j$, in which case we call $\theta$ the common Hermitian angle. As in the real case, we define $N^{\mathbb{C}}_{\alpha}(r)$ to be the maximum number of complex equiangular lines in $\mathbb{C}^{r}$ with common Hermitian angle $\arccos(\alpha)$ and $N^{\mathbb{C}}(r)=\max_{\alpha\in[0,1)}{N^{\mathbb{C}}_{\alpha}(r)}$.
The earliest results on complex equiangular lines go back to the 1975 work of Delsarte, Goethals, and Seidel [28]. Analogous to the real case, they proved the absolute bound
$N^{\mathbb{C}}(r)\leq r^{2}$
and gave matching lower bound constructions in $\mathbb{C}^{2}$ and $\mathbb{C}^{3}$. Zauner [91] was the first to make the connection between complex equiangular lines and quantum theory, as well as showing that $N^{\mathbb{C}}(r)=r^{2}$ for $r\leq 5$. This lead him to conjecture that the same holds for all $r\in\mathbb{N}$, and moreover that extremal constructions can be obtained as the orbit of some vector under the action of a Weyl–Heisenberg group.
Conjecture 1.10 (Zauner).
$N^{\mathbb{C}}(r)=r^{2}$ for all $r\in\mathbb{N}$.
Since the work of Renes, Blume-Kohout, Scott, and Caves [77], collections of $r^{2}$ complex equiangular lines in $\mathbb{C}^{r}$ have come to be known in quantum information theory as symmetric, informationally complete, positive operator-valued measures, or SIC-POVMs/SICs for short. Unlike the real case, Zauner’s conjecture has turned out to be true for more than 100 different values of the dimension $r$, including all $r\leq 40$ and as large as $r=19603$, see [47, 82, 4] and the survey paper of Fuchs, Hoang, and Stacey [43] for more information. In particular, Appleby, Bengtsson, Grassl, Harrison, and McConnell [4] provide a recipe which they conjecture produces a SIC in $\mathbb{C}^{r}$ for any $r\in\mathbb{N}$ such that $r-3$ is a perfect square. We also note that Szöllősi [83] classified all SICs in $\mathbb{C}^{3}$.
SICs turn out to be quite remarkable objects, having applications in quantum state tomography [21] and quantum cryptography [41], as well as being candidates for a “standard quantum measurement” in the foundations of quantum mechanics, most notably in QBism [42]. Moreover, they also have applications in high-precision radar and speech recognition and it has been suggested that a SIC in $\mathbb{C}^{2048}$ is worth patenting (see the last paragraph of [43]). Finally, it is very intriguing to note that Zauner’s conjecture is related to algebraic number theory with a connection to Hilbert’s twelfth problem [11, 5, 6, 4].
Despite all of the research on $N^{\mathbb{C}}(r)$, very little is known about $N^{\mathbb{C}}_{\alpha}(r)$. Delsarte, Goethels, and Seidel [28] proved the analogous relative bound
$N^{\mathbb{C}}_{\alpha}(r)\leq\frac{1-\alpha^{2}}{1-\alpha^{2}r}r\text{ for all }r<1/\alpha^{2},$
which is tight if and only if the unit vectors $v_{1},\ldots,v_{n}$ spanning the lines form a tight frame in $\mathbb{C}$, i.e. $\sum_{i=1}^{n}{v_{i}v_{i}^{*}}=\frac{n}{r}I$. To the best of our knowledge, no bound for general $\alpha$ is known for the complementary regime $r\geq 1/\alpha^{2}$. Since our projection method generalizes to the complex setting, we are able to obtain such bounds. Our results depends on a new geometric inequality for complex equiangular lines which is tight for any SIC and may therefore be of independent interest in quantum theory, see Remark 2.5.
The following theorem is a weaker, but more general version of Theorem 1.2, which generalizes the bound of Yu [90] and implies that $N^{\mathbb{C}}_{\alpha}(r)$ is linear in $r$ when $\alpha$ is fixed. Just like Theorem 1.2, the proof of Theorem 1.11 follows from bounds on the largest eigenvalue of a corresponding Gram matrix.
Theorem 1.11.
Let $10<\alpha<$ 1 and define $r^{\mathbb{C}}_{\alpha}=\frac{(1-\alpha)(1-\alpha^{2})}{\alpha^{3}}+3\frac{1+\alpha}{\alpha^{2}}$. For all $r\leq r^{\mathbb{C}}_{\alpha}$, we have
$N^{\mathbb{C}}_{\alpha}(r)\leq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$
with equality if and only if the corresponding lines form a SIC in $11/\alpha^{2}-$ 1 dimensions. Otherwise if $r>r^{\mathbb{C}}_{\alpha}$, then
$N^{\mathbb{C}}_{\alpha}(r)\leq\frac{1+\alpha}{\alpha}r+3\frac{1+\alpha}{\alpha^{3}}.$
Remark 1.12.
Similarly to the real case, it is well known (see e.g. [35]) that if there exists a SIC in $\mathbb{C}^{r}$, then $r=1/\alpha^{2}-1$, in which case Theorem 1.11 implies that for all $\frac{1}{\alpha^{2}}-1\leq r^{\prime}\leq r^{\mathbb{C}}_{\alpha}=\frac{1-O(\alpha)}{\alpha^{3}}$, we have $N^{\mathbb{C}}_{\alpha}(r^{\prime})=(1/\alpha^{2}-1)^{2}$. Conversely, if $N^{\mathbb{C}}_{\alpha}(r^{\prime})=(1/\alpha^{2}-1)^{2}$ for some $\frac{1}{\alpha^{2}}-1\leq r^{\prime}\leq r^{\mathbb{C}}_{\alpha}$, then Theorem 1.11 implies that $1/\alpha^{2}$ is an integer and the corresponding collection of lines forms a SIC in a $(1/\alpha^{2}-1)$-dimensional subspace.
In view of the fact that the existence of a SIC in $\mathbb{C}^{r}$ has been verified for many different $r$, we note that Remark 1.12 is more important than Remark 1.3. For instance, it shows that Zauner’s conjecture implies the following stronger conjecture.
Conjecture 1.13.
There exists a constant $C$ such that for all $10<\alpha<$ 1 satisfying $\frac{1}{\alpha^{2}}\in\mathbb{N}$ and for all $\frac{1}{\alpha^{2}}-1\leq r<\frac{1-C\alpha}{\alpha^{3}}$, we have
$N^{\mathbb{C}}_{\alpha}(r)=\left(\frac{1}{\alpha^{2}}-1\right)^{2}.$
We also note that Godsil and Roy [46] gave a construction of $r^{2}-r+1$ many equiangular lines in $\mathbb{C}^{r}$ with cosine of the common angle $\alpha=\sqrt{\frac{r-1}{r^{2}}}$ whenever $r-1$ is a prime power. This allows us to immediately obtain the following asymptotic variant of 1.13 as a corollary of Theorem 1.11.
Corollary 1.14.
Let $10<\alpha<$ 1 be such that $\frac{1+\sqrt{1-4\alpha^{2}}}{2\alpha^{2}}-1$ is a prime power and assume $\alpha\rightarrow 0$. Then for all $\frac{1+\sqrt{1-4\alpha^{2}}}{2\alpha^{2}}\leq r<\frac{1-o(1)}{\alpha^{3}}$, we have
$N^{\mathbb{C}}_{\alpha}(r)=\frac{1-o(1)}{\alpha^{4}}.$
We note that the construction of Godsil and Roy [46] is flat, i.e. the coordinates of the unit vectors which span the lines all have the same magnitude. They also showed that their construction is best possible, i.e. there are at most $r^{2}-r+1$ flat equiangular lines in $\mathbb{C}^{r}$. It would, therefore, be interesting to see if our methods can be used to obtain an exact version of Corollary 1.14 for flat equiangular lines.
1.3 Eigenvalues of regular graphs
Let $G$ be a $k$-regular graph whose adjacency matrix $A$ has second eigenvalue $\lambda_{2}$ and last eigenvalue $\lambda_{n}$. There is by now a lot of literature on how the size of $\lambda_{2}$ (and $\lambda_{n}$) relate to other properties of $G$, see [15, 25, 26] for more information. Inspired by the fact that $G$ is connected if and only if the spectral gap $k-\lambda_{2}>0$, Fiedler [36] studied the spectral gap as an algebraic measure of the connectivity of a graph and showed how it relates to the usual combinatorial notions of vertex and edge connectivity. It is also known that the spectral gap and $\max(\lambda_{2},-\lambda_{n})$ are related to the mixing time of random walks on a graph and that this generalises to other Markov chains, see e.g. [31, 32]. Moreover, Alon and Milman [1, 3] showed that the spectral gap as well as $\max(\lambda_{2},-\lambda_{n})$ are closely related to combinatorial notions of graph expansion, which can also be interpreted as discrete isoperimetric inequalities. In particular, Chung, Graham, and Wilson [22] showed that $\max(\lambda_{2},-\lambda_{n})$ determines how pseudo-random a graph is. We remark that the theory of graphs having good expansion properties is not only deep and interesting on its own, but also has extensive applications in mathematics and computer science [56, 65], as well as applications to other sciences, see e.g. [26, 85].
Given the above, it is interesting to determine how large the spectral gap $k-\lambda_{2}$ can be, or equivalently, how small $\lambda_{2}$ can be as a function of $k$. If $G$ has diameter $D$, the well-known Alon–Boppana theorem [1, 73] gives the lower bound $\lambda_{2}\geq 2\sqrt{k-1}-(2\sqrt{k-1}-1)/\lfloor D/2\rfloor$. As $D\rightarrow\infty$, this bound approaches $2\sqrt{k-1}$, which is known to be tight due to the existence of Ramanujan graphs, see e.g. [56]. Moreover, Alon conjectured and Friedman [39] proved that for any $\varepsilon>0$, “most” $k$-regular graphs on sufficiently many vertices will satisfy $\lambda_{2}\leq 2\sqrt{k-1}+\varepsilon$. However, the Alon–Boppana theorem doesn’t say much when the diameter $D\in\{2,3\}$ and even though this theorem has been improved and generalized in various ways (see e.g. [38, 55, 58]), it seems that all previously known results require the assumption that the graph is sufficiently sparse.
In this paper, we make use of the connection between graphs and equiangular lines in order to obtain the first generalization of the Alon–Boppana theorem which holds for dense $k$-regular graphs with no assumption on the diameter. We only require the following assumption on the spectral gap of $G$,
$k-\lambda_{2}<\frac{n}{2}.$ (1)
Indeed, such an assumption is necessary if we want to have a nontrivial lower bound on $\lambda_{2}$, since the complete bipartite graph with parts of size $n/2$ has $\lambda_{2}=0$. Moreover, note that any $k$-regular graph $G$ with $k<n/2$ must have $\lambda_{2}\geq 0$ (see Lemma 5.1), and so it must satisfy (1). Additionally, the complement of any $k$-regular graph is $(n-1-k)$-regular, so if $k\geq n/2$ then (1) holds for the complement graph. Finally, the bounds we obtain are tight for strongly regular graphs corresponding to families of equiangular lines meeting the absolute bound. To state the result, we let $\operatorname{m}_{G}(\lambda_{2})$ denote the multiplicity of $\lambda_{2}$ in $A$.
Theorem 1.15.
Let $k,n\in\mathbb{N}$ and let $G$ be a $k$-regular graph with second eigenvalue $\lambda_{2}=\lambda_{2}(G)$ and last eigenvalue $\lambda_{n}=\lambda_{n}(G)$. If the spectral gap satisfies $k-\lambda_{2}<n/2$, then we have
$2\left(k-\frac{(k-\lambda_{2})^{2}}{n}\right)\leq\frac{\lambda_{2}(\lambda_{2}+1)(2\lambda_{2}+1)}{1-\frac{2(k-\lambda_{2})}{n}}-\lambda_{2}(3\lambda_{2}+1),$
and
$-\lambda_{n}\leq\frac{\lambda_{2}(\lambda_{2}+1)}{1-\frac{2(k-\lambda_{2})}{n}}-\lambda_{2},$
with equality in both whenever $n=\binom{n-\operatorname{m}_{G}(\lambda_{2})+1}{2}-1$, i.e. when $G$ is a strongly regular graph corresponding to a family of $\binom{r+1}{2}$ equiangular lines in $\mathbb{R}^{r}$ for $r=n-\operatorname{m}_{G}(\lambda_{2})$.
Remark 1.16.
$k-\lambda_{2}<n/2$ implies that $k+\lambda_{2}<2\left(k-\frac{(k-\lambda_{2})^{2}}{n}\right)$, so that the first inequality of Theorem 1.15 gives an upper bound on the degree $k$.
Note that families of equiangular lines meeting the absolute bound $\binom{r+1}{2}$ in $\mathbb{R}^{r}$ are known to exist for $r=2,3,7,23$, and the corresponding graphs are the edge, the pentagon, the complement of the Schläfli graph, and the McLaughlin graph, so that the bounds of Theorem 1.15 are tight for such graphs. If we further assume that the spectral gap is bounded away from $n/2$, we immediately obtain the following corollary.
Corollary 1.17.
Let $k,n\in\mathbb{N}$ and let $G$ be a $k$-regular graph with second eigenvalue $\lambda_{2}=\lambda_{2}(G)$ and last eigenvalue $\lambda_{n}=\lambda_{n}(G)$. If the spectral gap satisfies $k-\lambda_{2}\leq(1-\varepsilon)\frac{n}{2}$ for some constant $\varepsilon>0$, then we have
[…]
In particular, if $G$ is bipartite then $\lambda_{2}\geq\Omega(\sqrt{k}).$ Moreover, if the spectral gap $k-\lambda_{2}=o(n)$ then
[…]
so that if $G$ is bipartite, we have $\lambda_{2}\geq(1-o(1))\sqrt{k}$.
As mentioned in Remark 1.1, Räty, Sudakov, and Tomon [76] noted that R. Metz (see [16], Section 7.A) constructed an infinite family of strongly regular graphs satisfying $k=(1-o(1))(2n)^{3/4}$, $\lambda_{2}=(1-o(1))k^{1/3}$, and $-\lambda_{n}=\lambda_{2}^{2}$ and moreover, that for a fixed odd prime $p$, Davis, Huczynska, Johnson, and Polhill [27] recently constructed an infinite family of strongly regular graphs satisfying $k=(1-o(1))n/p$, $\lambda_{2}=(1-o(1))(pk)^{1/3}$, and $-\lambda_{n}=(1-o(1))\lambda_{2}^{2}/p$. It follows that Corollary 1.17 is tight when $k=(1-o(1))(2n)^{3/4}$ or $k=\Theta(n)$.
Remark 1.18.
Note that for any prime power $q$, if we let $\alpha=1/(2q-1)$, then the aformentioned construction of Metz leads to a collection of $\Theta(1/\alpha^{4})$ equiangular lines with common angle $\arccos(\alpha)$ in $\mathbb{R}^{r}$ with $r=\Theta(1/\alpha^{3})$, matching the upper bound of Theorem 1.2.
1.4 Notation and definitions
Let $\mathbb{N},\mathbb{R},\mathbb{C}$ denote the set of positive integers, field of real numbers, and field of complex numbers, respectively. For all $n\in\mathbb{N}$, we define $[n]=\{1,\ldots,n\}$. We let $S^{1}(\alpha)=\{z\in\mathbb{C}:|z|=\alpha\}$ denote the circle of radius $\alpha$ centered at 00 in the complex plane. For any quantities $f,g$ we write $f\leq O(g)$ or $g\geq\Omega(f)$, if $f,g$ are implicitly assumed to be functions of an increasing parameter $t\in\mathbb{N}$ and there exist $C>0,t_{0}\in\mathbb{N}$ such that $f(t)\leq Cg(t)$ for all $t\geq t_{0}$. Moreover, we write $f=\Theta(g)$ if $f\leq O(g)$ and $f\geq\Omega(g)$. Also, we write $f=o(g)$, $g=\omega(f)$ or $f\ll g$ whenever $f(t)/g(t)\rightarrow 0$ as $t\rightarrow\infty$.
Linear algebra
Let $\mathbb{F}$ be a field and let $S$ be a finite set. We let $\mathbb{F}^{S}$ denote the vector space of functions from $S$ to $\mathbb{F}$, we let $\mathbbm{1}$ denote the all ones vector (i.e. the constant function with value $1$), and for $i\in S$, we let $e_{i}$ denote the standard basis vector in $\mathbb{F}^{S}$ given by $(e_{i})(j)=1$ if $i=j$ and $(e_{i})(j)=0$ otherwise. More generally, given another finite set $T$, we shall consider functions in $\mathbb{F}^{S\times T}$ as matrices indexed by $S\times T$ and we also do no distinguish between matrices in $\mathbb{F}^{S\times T}$ and the corresponding linear maps from $\mathbb{F}^{T}$ to $\mathbb{F}^{S}$. For a matrix $M\in\mathbb{F}^{S\times S}$, we let $\operatorname{rk}(M)$ and $\operatorname{tr}(M)$ denote its rank and trace, respectively and for a pair of matrices $M,N\in\mathbb{F}^{S\times S}$, we let $M\odot N$ denote their Hadamard product, given by $(M\odot N)(i,j)=M(i,j)N(i,j)$.
Given $M\in\mathbb{C}^{S\times T}$, we let $M^{\intercal}$ and $M^{*}$ denote the transpose matrix and adjoint matrix in $\mathbb{C}^{T\times S}$. We also let $\mathscr{S}_{S}=\{M\in\mathbb{R}^{S\times S}:M^{\intercal}=M\}$ denote the space of all real symmetric matrices indexed by $S\times S$ and let $\left\langle u,v\right\rangle=u^{*}v$ denote the standard inner product of $u,v\in\mathbb{C}^{S}$. For any Hermitian matrix $M\in\mathbb{C}^{S\times S}$, we let $\lambda_{i}=\lambda_{i}(M)\in\mathbb{R}$ for $i\in S$ denote its eigenvalues ordered so that $\lambda_{1}\geq\ldots\geq\lambda_{n}$ and we define $\operatorname{m}_{M}(\lambda)=|\{j:\lambda_{j}=\lambda\}|$ to be the multiplicity of $\lambda$. Finally, letting $u_{1},\ldots,u_{n}$ be an orthonormal basis of corresponding eigenvectors such that $M=\sum_{i=1}^{n}{\lambda_{i}u_{i}u_{i}^{*}}$, we define the Moore–Penrose inverse of $M$ to be $M^{{\dagger}}=\sum_{i:\lambda_{i}\neq 0}{\frac{1}{\lambda_{i}}u_{i}u_{i}^{*}}$.
For convenience, we let $\mathbb{F}^{n}=\mathbb{F}^{[n]}$, $\mathbb{F}^{n\times S}=\mathbb{F}^{[n]\times S}$, and $\mathscr{S}_{n}=\mathscr{S}_{[n]}$.
Inner product spaces
Let $\mathbb{F}\in\{\mathbb{R},\mathbb{C}\}$ and let $V$ be a vector space over $\mathbb{F}$. A map $\left\langle\cdot,\cdot\right\rangle_{\text{P}}\colon V\times V\rightarrow\mathbb{F}$ is called an inner product if it satisfies
  1. 1.
    $\left\langle u,\cdot\right\rangle_{\text{P}}\colon V\rightarrow\mathbb{F}$ is linear for all $u\in V$,
  2. 2.
    $\overline{\left\langle u,v\right\rangle_{\text{P}}}=\left\langle v,u\right\rangle_{V}$ for all $u,v\in V$,
  3. 3.
    $\left\langle u,u\right\rangle_{\text{P}}\ =0$ implies $u=0$ for all $u\in V$,
and we say that $V$ is an inner product space with respect to $\left\langle\cdot,\cdot\right\rangle_{\text{P}}$, letting $||\cdot||_{\text{P}}\colon V\rightarrow\mathbb{F}$ denote the corresponding norm. For any $u\in V$, we let $u^{\#}\colon V\rightarrow\mathbb{F}$ denote the adjoint of $u$ defined by $u^{\#}v=\left\langle u,v\right\rangle_{\text{P}}$ for all $v\in V$. More generally, if $U$ and $V$ are finite-dimensional vector spaces with inner products $\left\langle\cdot,\cdot\right\rangle_{\text{P}}$ and $\left\langle\cdot,\cdot\right\rangle_{\text{Q}}$, respectively and $L\colon U\rightarrow V$ is a linear map, then we let $L^{\#}\colon V\rightarrow U$ denote the adjoint map [^5] If we choose bases for $U$ and $V$, then we can represent the linear map $L$ by a corresponding matrix $[L]$ with respect to those bases, in which case the adjoint map $L^{\#}$ is represented by the corresponding adjoint matrix $[L]^{*}$., which satisfies $\left\langle L^{\#}v,u\right\rangle_{\text{P}}=\left\langle v,Lu\right\rangle_{\text{Q}}$ for all $u\in U,v\in V$.
Let $V$ be an inner product space with respect to $\left\langle\cdot,\cdot\right\rangle_{\text{P}}$ and let $\mathscr{C}\subseteq V$ be a finite set of vectors. We define $M_{\mathscr{C}}\in\mathbb{F}^{\mathscr{C}\times\mathscr{C}}$ to be their Gram matrix, defined by $(M_{\mathscr{C}})(u,v)=\left\langle u,v\right\rangle_{\text{P}}$. Moreover, for any subset $S\subseteq\mathbb{F}$, we say that a collection of unit vectors $\mathscr{C}$ is a spherical $S$-code if $\left\langle u,v\right\rangle_{\text{P}}\in S$ for all $u\neq v\in\mathscr{C}$. Finally, given a finite set $S$, we let $\left\langle\cdot,\cdot\right\rangle_{F}\colon\mathbb{F}^{S\times S}\times\mathbb{F}^{S\times S}\rightarrow\mathbb{F}$ denote the Frobenius inner product, which is defined by $\left\langle A,B\right\rangle_{F}=\operatorname{tr}\left(A^{*}B\right)$ for all $A,B\in\mathbb{F}^{S\times S}$.
Graph theory
A graph $G$ consists of a finite set $\operatorname{V}(G)$ whose elements are called vertices and a collection of unordered pairs of vertices $\operatorname{E}(G)\subseteq\{\{u,v\}:u,v\in\operatorname{V}(G),u\neq v\}$, which are called edges. For convenience, whenever $u$ and $v$ are vertices of a graph $G$, we write $uv$ instead of $\{u,v\}$ in order to represent the unordered pair. We let $G^{c}$ denote the complement of $G$ and we write $H\subseteq G$ when $H$ is a subgraph of $G$. For any vertex $v\in V(G)$, we let $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ denote its neighborhood and let $d_{G}(v)=|N_{G}(v)|$ denote its degree. We also define $\overline{d_{G}}=\frac{1}{|V(G)|}\sum_{v\in V(G)}{d_{G}(v)}$ to be the average degree of $G$ and $\Delta_{G}=\max_{v\in V(G)}{d_{G}(v)}$ to be the maximum degree of $G$. A graph $G$ is called $d$-regular if $d_{G}(v)=d$ for all $v\in V(G)$. We let $K_{n}$ denote the complete graph on $n$ vertices and for $r\geq 2$, let $K_{n_{1},\ldots,n_{r}}$ denote the complete multipartite graph with $r$ parts having sizes $n_{1},\ldots,n_{r}$.
For any graph $G$, we let $A_{G}\in\mathbb{R}^{V(G)\times V(G)}$ denote the corresponding adjacency matrix, defined by $A_{G}(u,v)=1$ if $uv\in E(G)$ and $A_{G}(u,v)=0$ otherwise. Since $A_{G}$ is symmetric, it has a full set of real eigenvalues $\lambda_{1}(G)\geq\ldots\geq\lambda_{n}(G)$, and we equivalently say that $\lambda_{i}=\lambda_{i}(G)$ is the $i$th eigenvalue of $G$ with multiplicity $\operatorname{m}_{G}(\lambda_{i})=\operatorname{m}_{A_{G}}(\lambda_{i})$. In particular, $\lambda_{1}$ is called the spectral radius of $G$ and we say that $\lambda_{1}-\lambda_{2}$ is the spectral gap of $G$. Finally, the spectral radius order $k(\lambda)$ is defined to be the least number of vertices in a graph having spectral radius $\lambda$, where we set $k(\lambda)=\infty$ if no such graph exists.
2 Orthogonal projection of matrices
In this section, we discuss how orthogonal projection of matrices with respect to the Frobenius inner product can be used to obtain new inequalities for positive semidefinite matrices (equivalently Gram matrices of vectors) which underlie our main results. We also apply this method to obtain a strengthening of the first Welch bound in coding theory, in particular giving another proof of the relative bound. Finally, we discuss how our approach relates to the Delsarte linear programming method.
Real equiangular lines
The usual way one studies real equiangular lines in $\mathbb{R}^{r}$ is by choosing a unit vector along each line, thereby obtaining a collection of vectors $\mathscr{C}$ with all pairwise inner products lying in the set $\{\alpha,-\alpha\}$ where $\alpha$ is the cosine of the common angle, i.e. $\mathscr{C}$ is a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$. In our previous work together with Dräxler, Keevash, and Sudakov [7], the arguments rested upon finding a sufficiently large subset of $\mathscr{C}$ with the property that all pairwise inner products are $\alpha$, i.e. a large regular simplex, and then projecting remaining vectors onto the orthogonal complement of the span of this simplex. However, instead of looking for such a simplex in $\mathscr{C}$, in this paper we use the fact that the projection matrices $\{vv^{\intercal}:v\in\mathscr{C}\}$ already form a large regular simplex in $\mathscr{S}_{r}$, the space of $r\times r$ real symmetric matrices with respect to the Frobenius inner product. This result is well known and indeed, when combined with the fact that the dimension of $\mathscr{S}_{r}$ is $\binom{r+1}{2}$, it yields a proof of the absolute bound.
Our main contribution is the observation that one can obtain new geometric inequalities by orthogonally projecting a matrix onto the span of $\{vv^{\intercal}:v\in\mathscr{C}\}$ and using the fact that the Frobenius norm can only decrease. A benefit of this approach is that if the absolute bound is met, i.e. if $n=\binom{r+1}{2}$, then $\{vv^{\intercal}:v\in\mathscr{C}\}$ span the whole space $\mathscr{S}_{r}$, and so our inequalities become tight.
So what kind of matrices should we project onto $\{vv^{\intercal}:v\in\mathscr{C}\}$? Since an orthogonal projection of an asymmetric matrix onto a space of symmetric matrices necessarily reduces its Frobenius norm and since we are interested in obtaining inequalities that have the possibility of being tight, we restrict ourselves to projecting symmetric matrices. Perhaps the simplest matrix to try is the identity and indeed, this leads to a proof of the relative bound (see Remark 2.7). It is also natural to try to construct the matrix as a tensor product of some specially chosen vectors. In particular, in this paper we observe that projecting $\left(\sum_{v\in\mathscr{C}}{x(v)v}\right)u^{\intercal}+u\left(\sum_{v\in\mathscr{C}}{x(v)v}\right)^{\intercal}$, where $u\in\mathscr{C}$ and $x$ is a unit eigenvector corresponding to eigenvalue $\lambda$ of the Gram matrix $M_{\mathscr{C}}$, leads to the following inequality involving $\alpha,n,x(u)$, and $\lambda$, which is tight if $n=\binom{r+1}{2}$.
Theorem 2.1.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with Gram matrix $M=M_{\mathscr{C}}$ and $n=|\mathscr{C}|$. If $\lambda$ is an eigenvalue of $M$ with corresponding unit eigenvector $x$, then
$\lambda\left(\frac{\lambda^{2}}{\alpha^{2}n+1-\alpha^{2}}-\frac{1-\alpha^{2}}{2\alpha^{2}}\right)x(u)^{2}\geq\lambda-\frac{1-\alpha^{2}}{2\alpha^{2}}$
for all $u\in\mathscr{C}$, with equality if $n=\binom{r+1}{2}$.
For the proof of Theorem 2.1, recall that given a linear map $\mathscr{L}\colon\mathbb{R}^{\mathscr{C}}\rightarrow\mathscr{S}_{r}$ where $\mathscr{S}_{r}$ is equipped with the Frobenius inner product and $\mathbb{R}^{\mathscr{C}}$ is equipped with the standard inner product, the adjoint #$\mathscr{L}^{\#}\colon\mathscr{S}_{r}\rightarrow\mathbb{R}^{n}$ satisfies #$\left\langle\mathscr{L}^{\#}M,v\right\rangle=\left\langle M,\mathscr{L}v\right\rangle_{F}$ and for a single matrix $M\in\mathscr{S}_{r}$, the adjoint $M^{\#}\colon\mathscr{S}_{r}\rightarrow\mathbb{R}$ satisfies $M^{\#}N=\left\langle M,N\right\rangle_{F}$. We will also make use of the well-known fact that $\left\langle xy^{\intercal},zw^{\intercal}\right\rangle_{F}=\left\langle x,z\right\rangle\left\langle y,w\right\rangle$ for any $x,y,z,w\in\mathbb{R}^{r}$.
Remark 2.2.
If we let $\mathscr{C}=\{v_{1},\ldots,v_{n}\}$, then since any $r\times r$ matrix $A$ can be viewed as a vector $\operatorname{vec}(A)$ with $r^{2}$ coordinates, one can equivalently think of $\mathscr{W}$ in the following proof as an $r^{2}\times n$ matrix $W$ with $i$th column given by $\operatorname{vec}(v_{i}v_{i}^{\intercal})$. In this case, the Frobenius inner product simply becomes the standard inner product in $\mathbb{R}^{r^{2}}$ and the adjoint map $\mathscr{W}^{\#}$ becomes the $n\times r^{2}$ transpose matrix $W^{\intercal}$.
Proof of Theorem 2.1.
We define $\mathscr{W}\colon\mathbb{R}^{\mathscr{C}}\rightarrow\mathscr{S}_{r}$ to be the linear map given by $\mathscr{W}e_{v}=vv^{\intercal}$ for all $v\in\mathscr{C}$ and consider the adjoint map $\mathscr{W}^{\#}\colon\mathscr{S}_{r}\rightarrow\mathbb{R}^{\mathscr{C}}$ with respect to the Frobenius inner product on $\mathscr{S}_{r}$ and the standard inner product on $\mathbb{R}^{\mathscr{C}}$. Observe that $\mathscr{W}^{\#}\mathscr{W}\colon\mathbb{R}^{\mathscr{C}}\rightarrow\mathbb{R}^{\mathscr{C}}$ is a matrix satisfying
$(\mathscr{W}^{\#}\mathscr{W})(u,v)=\left\langle uu^{\intercal},vv^{\intercal}\right\rangle_{F}=\left\langle u,v\right\rangle^{2}=\begin{cases}\alpha^{2}&\text{ if }u\neq v\\ 1&\text{ if }u=v\end{cases}$
for all $u,v\in\mathscr{C}$, so that $\mathscr{W}^{\#}\mathscr{W}=(1-\alpha^{2})I+\alpha^{2}J$, which is invertible with $(\mathscr{W}^{\#}\mathscr{W})^{-1}=\frac{1}{1-\alpha^{2}}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)$. It follows that $\mathscr{P}=\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}\colon\mathscr{S}_{r}\rightarrow\mathscr{S}_{r}$ is the orthogonal projection onto the range of $\mathscr{W}$. Indeed, one may verify that $\mathscr{P}^{2}=\mathscr{P}=\mathscr{P}^{\#}$ and $\mathscr{P}\mathscr{W}=\mathscr{W}$.
Now let $u\in\mathscr{C}$, let $V\in\mathbb{R}^{r\times\mathscr{C}}$ be the matrix satisfying $Vy=\sum_{v\in\mathscr{C}}{y(v)v}$ for all $y\in\mathbb{R}^{\mathscr{C}}$, and let $X=\left(Vx\right)u^{\intercal}+u\left(Vx\right)^{\intercal}$. Since $V^{\intercal}V=M$, we have
$||X||_{F}^{2}=\left\langle(Vx)u^{\intercal}+u(Vx)^{\intercal},(Vx)u^{\intercal}+u(Vx)^{\intercal}\right\rangle_{F}$ $=2\left\langle Vx,Vx\right\rangle\left\langle u,u\right\rangle+2\left\langle Vx,u\right\rangle^{2}$
$=2\left(x^{\intercal}Mx+(x^{\intercal}Me_{u})^{2}\right)$
$=2\left(\lambda+\lambda^{2}x(u)^{2}\right).$
Moreover, for all $v\in\mathscr{C}$ we have
$(\mathscr{W}^{\#}X)(v)=\left\langle vv^{\intercal},(Vx)u^{\intercal}+u(Vx)^{\intercal}\right\rangle_{F}=2\left\langle v,Vx\right\rangle\left\langle v,u\right\rangle=2(e_{v}^{\intercal}Mx)\left\langle v,u\right\rangle=2\lambda x(v)M(v,u),$
so that $\left\langle\mathbbm{1},\mathscr{W}^{\#}X\right\rangle=2\lambda\sum_{v\in\mathscr{C}}{x(v)M(v,u)}=2\lambda x^{\intercal}Me_{u}=2\lambda^{2}x(u)$ and
$||\mathscr{W}^{\#}X||^{2}=4\lambda^{2}\sum_{v\in\mathscr{C}}{x(v)^{2}\left\langle v,u\right\rangle^{2}}$ $=4\lambda^{2}\left(\alpha^{2}\sum_{v\in\mathscr{C}}{x(v)^{2}}+(1-\alpha^{2})x(u)^{2}\right)$
$=4\lambda^{2}\left(\alpha^{2}+(1-\alpha^{2})x(u)^{2}\right).$
Since $\mathscr{P}$ is an orthogonal projection, we conclude that
$2\left(\lambda+\lambda^{2}x(u)^{2}\right)=||X||^{2}_{F}\geq||\mathscr{P}X||_{F}^{2}$ $=X^{\#}\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}X$
$=\frac{1}{1-\alpha^{2}}(\mathscr{W}^{\#}X)^{\intercal}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)(\mathscr{W}^{\#}X)$
$=\frac{1}{1-\alpha^{2}}\left(||\mathscr{W}^{\#}X||^{2}-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}\left\langle\mathbbm{1},\mathscr{W}^{\#}X\right\rangle^{2}\right)$
$=\frac{1}{1-\alpha^{2}}\left(4\lambda^{2}\left(\alpha^{2}+(1-\alpha^{2})x(u)^{2}\right)-\frac{4\alpha^{2}\lambda^{4}x(u)^{2}}{\alpha^{2}n+1-\alpha^{2}}\right),$
which is equivalent to the desired inequality.
Finally, since $\mathscr{W}^{\#}\mathscr{W}$ is invertible we have that $n=\operatorname{rk}(\mathscr{W}^{\#}\mathscr{W})=\operatorname{rk}(\mathscr{W})$. Also note that the space of symmetric $r\times r$ matrices $\mathscr{S}_{r}$ has dimension $\binom{r+1}{2}$. Thus if $n=\binom{r+1}{2}$, then the range of $\mathscr{W}$ is $\mathscr{S}_{r}$ and so $\mathscr{P}$ is the identity map, giving equality above. ∎
Remark 2.3.
We note that the preceding theorem can also be obtained via an application of the Schur product theorem. Indeed, the projection inequality in the above is equivalent to $||X-\sum_{v\in\mathscr{C}}{c(v)vv^{\intercal}}||^{2}_{F}\geq 0$ where $X=(Vx)u^{\intercal}+u(Vx)^{\intercal}$ and the coefficients $\{c(v):v\in\mathscr{C}\}$ are given by the projection of $X$ onto the span of $\{vv^{\intercal}:v\in\mathscr{C}\}$. Since $X=(Vx+u)(Vx+u)^{\intercal}-(Vx)(Vx)^{\intercal}-uu^{\intercal}$, this inequality can be written in the form $\left|\left|\sum_{v\in\mathscr{C}^{\prime}}{c(v)vv^{\intercal}}\right|\right|^{2}_{F}\geq 0$ where $\mathscr{C}^{\prime}=\mathscr{C}\cup\{Vx,u,Vx+u\}$ and $c(Vx)=c(u)=-1,c(Vx+u)=1$. Now it suffices to observe that $\left|\left|\sum_{v\in\mathscr{C}^{\prime}}{c(v)vv^{\intercal}}\right|\right|^{2}_{F}=\sum_{u,v\in\mathscr{C}^{\prime}}{c(u)c(v)\left\langle u,v\right\rangle^{2}}=c^{\intercal}(M\odot M)c$ where $M=M_{\mathscr{C}^{\prime}}$ is the Gram matrix of $\mathscr{C}^{\prime}$ and $\odot$ is the Hadamard product. Therefore, the desired inequality is equivalent to $c^{\intercal}(M\odot M)c\geq 0$, which follows from the Schur product theorem applied to $M$.
Complex equiangular lines
Another benefit of our approach is that it generalizes to the setting of complex equiangular lines, in which case we have a collection of unit vectors $\mathscr{C}$ in $\mathbb{C}^{r}$ satisfying $|\left\langle u,v\right\rangle|=\alpha$ for all $u\neq v\in\mathscr{C}$, i.e. a spherical $S^{1}(\alpha)$-code where $S^{1}(\alpha)$ is the circle of radius $\alpha$ centered at 00 in the complex plane. As in the real case, it is well known that the Hermitian matrices $\{vv^{*}:v\in\mathscr{C}\}$ form a regular simplex with respect to the Frobenius inner product, so that we have the analogous absolute bound $|\mathscr{C}|\leq r^{2}$. However, unlike in the real case, a complex linear combination of Hermitian matrices need not be Hermitian. Indeed, even for a single Hermitian matrix $H$ we have $(iH)^{*}=-iH\neq iH$. Therefore, $\{vv^{*}:v\in\mathscr{C}\}$ is capable of spanning the entire space $\mathbb{C}^{r\times r}$ of $r\times r$ complex matrices and so we do not need to restrict ourselves to projecting Hermitian matrices. As such, if we let $u\in\mathscr{C}$ and let $x$ be a unit eigenvector of the corresponding Gram matrix $M_{\mathscr{C}}$, then orthogonally projecting $\left(\sum_{v\in\mathscr{C}}{x(v)v}\right)u^{*}$ onto the span of $\{vv^{*}:v\in\mathscr{C}\}$ yields the following result which is analogous to Theorem 2.1.
Theorem 2.4.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $S^{1}(\alpha)$-code in $\mathbb{C}^{r}$ with Gram matrix $M=M_{\mathscr{C}}$ and $n=|\mathscr{C}|$. If $\lambda$ is an eigenvalue of $M$ with corresponding unit eigenvector $x$, then
$\lambda\left(\frac{\lambda^{2}}{\alpha^{2}n+1-\alpha^{2}}-\frac{1-\alpha^{2}}{\alpha^{2}}\right)|x(u)|^{2}\geq\lambda-\frac{1-\alpha^{2}}{\alpha^{2}}$
for all $u\in\mathscr{C}$, with equality if $n=r^{2}$.
The proof of Theorem 2.4 is almost identical to (and arguably simpler than) that of Theorem 2.1, so we omit some of the details. For the reader who is interested in filling in these details, it will be helpful to note that $V^{*}V=M$ and $\left\langle xy^{*},zw^{*}\right\rangle_{F}=\left\langle x,z\right\rangle\overline{\left\langle y,w\right\rangle}$ for all $x,y,z,w\in\mathbb{C}^{r}$.
Proof of Theorem 2.4.
Define the linear map $\mathscr{W}\colon\mathbb{C}^{\mathscr{C}}\rightarrow\mathbb{C}^{r\times r}$ by $\mathscr{W}e_{v}=vv^{*}$, so that we have $(\mathscr{W}^{\#}\mathscr{W})^{-1}=\frac{1}{1-\alpha^{2}}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)$ and let $\mathscr{P}=\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}$ be the corresponding orthogonal projection map. Also let $u\in\mathscr{C}$, let $V\in\mathbb{C}^{r\times\mathscr{C}}$ be the matrix satisfying $Vy=\sum_{v\in\mathscr{C}}{y(v)v}$ for all $y\in\mathbb{C}^{\mathscr{C}}$, and let $X=(Vx)u^{*}$. Then we have
$\lambda=||X||_{F}\geq||\mathscr{P}X||_{F}$ $=X^{\#}\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}X$
$=\frac{1}{1-\alpha^{2}}(\mathscr{W}^{\#}X)^{\intercal}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)(\mathscr{W}^{\#}X)$
$=\frac{1}{1-\alpha^{2}}\left(||\mathscr{W}^{\#}X||^{2}_{F}-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}\left\langle\mathbbm{1},\mathscr{W}^{\#}X\right\rangle^{2}\right)$
$=\frac{1}{1-\alpha^{2}}\left(\lambda^{2}\left(\alpha^{2}+(1-\alpha^{2})|x(u)|^{2}\right)-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}\lambda^{4}|x(u)|^{2}\right),$
which is equivalent to the desired bound. Moreover, if $n=r^{2}$ then $\{vv^{*}:v\in\mathscr{C}\}$ span $\mathbb{C}^{r\times r}$ and so $\mathscr{P}$ is the identity map, giving equality above. ∎
Remark 2.5.
Recalling that $\odot$ denotes the Hadamard product, we note that projecting $\left(Vx\right)\left(Vy\right)^{*}$ onto the span of $\{vv^{*}:v\in\mathscr{C}\}$ yields the inequality
$(1-\alpha^{2})\left\langle x,Mx\right\rangle\left\langle y,My\right\rangle+\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}|\left\langle Mx,My\right\rangle|^{2}\geq\left\langle Mx\odot\overline{Mx},My\odot\overline{My}\right\rangle,$
which is tight if $n=r^{2}$. Moreover, the case of equality in the above could be of particular interest because sets of $r^{2}$ equiangular lines in $\mathbb{C}^{r}$ are important objects in quantum information theory known as SICs, and 1.10 implies that they exist for every $r$, see Section 1.2 for more information. In particular, it is well known that a set of $r^{2}$ equiangular lines in $\mathbb{C}^{r}$ with common angle $\arccos(\alpha)$ must satisfy $\alpha=1/\sqrt{r+1}$ (see e.g. [28]). Therefore, for all vectors $x,y\in\mathbb{R}^{\mathscr{C}}$, the Gram matrix $M$ of a SIC satisfies
$r\left\langle x,Mx\right\rangle\left\langle y,My\right\rangle+\frac{1}{r}|\left\langle Mx,My\right\rangle|^{2}=(r+1)\left\langle Mx\odot\overline{Mx},My\odot\overline{My}\right\rangle.$
A refinement of the first Welch bound
It is well known that the relative bound is actually a special case of the first Welch bound, an inequality in coding theory [89] with applications in signal analysis and quantum information theory, see [34, 68, 70]. In a more general form, this bound states that for any collection $\mathscr{C}$ of $n$ unit vectors in $\mathbb{C}^{r}$, we have
$\sum_{u,v\in\mathscr{C}}{|\left\langle u,v\right\rangle|^{2}}\geq\frac{n^{2}}{r}.$
Moreover, the first Welch bound follows from the inequality $\left|\left|I-\sum_{v\in\mathscr{C}}{c(v)vv^{*}}\right|\right|_{F}^{2}\geq 0$ where $c(v)=1/n$ for all $v\in\mathscr{C}$, and so choosing coefficients $\{c(v):v\in\mathscr{C}\}$ such that $||I-\sum_{v\in\mathscr{C}}{c(v)vv^{*}}||_{F}^{2}$ is minimized would clearly strengthen the bound. Since choosing such coefficients is precisely equivalent to orthogonally projecting the identity $I$ onto the span of $v_{1}v_{1}^{*},\ldots,v_{n}v_{n}^{*}$, our method allows us to obtain a refinement of the first Welch bound.
To state our result, recall that for any Hermitian matrix $M$ with eigenvalues $\lambda_{1},\ldots,\lambda_{n}$ and corresponding unit eigenvectors $u_{1},\ldots,u_{n}$, the Moore–Penrose generalized inverse of $M$ is $M^{{\dagger}}=\sum_{i:\lambda_{i}>0}{\frac{1}{\lambda_{i}}u_{i}u_{i}^{\intercal}}$. Also note that $\odot$ denotes the Hadamard product so that the entries of $M\odot\overline{M}$ are given by $\left(M\odot\overline{M}\right)(i,j)=|M(i,j)|^{2}$.
Theorem 2.6.
Let $n\in\mathbb{N}$ and let $\mathscr{C}$ be a collection of $n$ unit vectors in $\mathbb{C}^{r}$ with corresponding Gram matrix $M=M_{\mathscr{C}}$. Then
$\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{{\dagger}}\mathbbm{1}\leq r,$
with equality if and only if the $r\times r$ identity matrix is in the span of $\{vv^{*}:v\in\mathscr{C}\}$. Moreover
$\sum_{u,v\in\mathscr{C}}{|\left\langle u,v\right\rangle|^{2}}\geq\frac{n^{2}}{r}\left(2-\frac{\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{{\dagger}}\mathbbm{1}}{r}\right),$
and when $M\odot\overline{M}$ is invertible, we have equality if and only if $\left(M\odot\overline{M}\right)\mathbbm{1}=\frac{n}{r}\mathbbm{1}$.
Proof.
Define the linear map $\mathscr{W}\colon\mathbb{C}^{\mathscr{C}}\rightarrow\mathbb{C}^{r\times r}$ by $We_{v}=vv^{*}$ and consider the adjoint map $\mathscr{W}^{\#}\colon\mathbb{C}^{r\times r}\rightarrow\mathbb{C}^{\mathscr{C}}$. We have that
$(\mathscr{W}^{\#}\mathscr{W})(u,v)=\left\langle u,v\right\rangle\overline{\left\langle u,v\right\rangle}=|\left\langle u,v\right\rangle|^{2}$
for all $u,v\in\mathscr{C}$, so that $\mathscr{W}^{\#}\mathscr{W}=M\odot\overline{M}$. Observe that the $r\times r$ identity matrix $I$ satisfies $\mathscr{W}^{\#}I=\mathbbm{1}$. Moreover, it is well known (see e.g. [12]) that $\mathscr{P}=\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{{\dagger}}\mathscr{W}^{\#}\colon\mathbb{C}^{r\times r}\rightarrow\mathbb{C}^{r\times r}$ is the orthogonal projection onto the range of $\mathscr{W}$, so that we obtain our first inequality
$r=||I||^{2}_{F}\geq||\mathscr{P}I||_{F}^{2}=\left\langle I,\mathscr{P}I\right\rangle_{F}=I^{\#}\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{{\dagger}}\mathscr{W}^{\#}I=\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{{\dagger}}\mathbbm{1}.$
Moreover, because $\mathscr{P}$ is a projection, we have equality above if and only if $\mathscr{P}I=I$, which occurs if and only if $I$ lies in the span of $\{vv^{*}:v\in\mathscr{C}\}$.
Now we use the fact that $\mathscr{P}I$ is the unique minimizer of the expression $||I-X||_{F}^{2}$ over all $X$ in the range of $\mathscr{W}$, in order to conclude that
$r-\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{\dagger}\mathbbm{1}=\left|\left|I-\mathscr{P}I\right|\right|_{F}^{2}\leq\left|\left|I-\frac{r}{n}\mathscr{W}\mathbbm{1}\right|\right|_{F}^{2}$ $=I^{\#}I-\frac{2r}{n}I^{\#}\mathscr{W}\mathbbm{1}+\frac{r^{2}}{n^{2}}\mathbbm{1}^{\intercal}\mathscr{W}^{\#}\mathscr{W}\mathbbm{1}$
$=r-\frac{2r}{n}\mathbbm{1}^{\intercal}\mathbbm{1}+\frac{r^{2}}{n^{2}}\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)\mathbbm{1}$
$=\frac{r^{2}}{n^{2}}\sum_{u,v\in\mathscr{C}}{|\left\langle u,v\right\rangle|^{2}}-r,$
which is equivalent to the second desired inequality.
Furthermore, let us assume that $M\odot\overline{M}$ is invertible, so that we have $\mathscr{P}I=\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{{\dagger}}\mathscr{W}^{\#}I=\mathscr{W}\left(M\odot\overline{M}\right)^{-1}\mathbbm{1}$. By the uniqueness of $\mathscr{P}I$, the previous inequality is tight if and only if $\mathscr{P}I=\frac{r}{n}\mathscr{W}\mathbbm{1}$. Thus if $\mathscr{P}I=\frac{r}{n}\mathscr{W}\mathbbm{1}$, then we conclude
$\frac{r}{n}\left(M\odot\overline{M}\right)\mathbbm{1}=\frac{r}{n}\mathscr{W}^{\#}\mathscr{W}\mathbbm{1}=\mathscr{W}^{\#}\mathscr{P}I=\mathscr{W}^{\#}\mathscr{W}\left(\mathscr{W}^{\#}\mathscr{W}\right)^{-1}\mathscr{W}^{\#}I=\mathbbm{1}.$
On the other hand, if $\left(M\odot\overline{M}\right)\mathbbm{1}=\frac{n}{r}\mathbbm{1}$ then $\left(M\odot\overline{M}\right)^{-1}\mathbbm{1}=\frac{r}{n}\mathbbm{1}$ and applying $\mathscr{W}$ yields $\mathscr{P}I=\mathscr{W}\left(\mathscr{W}^{\#}\mathscr{W}\right)^{-1}\mathscr{W}^{\#}I=\mathscr{W}\left(M\odot\overline{M}\right)^{-1}\mathbbm{1}=\frac{r}{n}\mathscr{W}\mathbbm{1}$, as desired. ∎
Remark 2.7.
Note that if $\mathscr{C}$ is a spherical $\{\alpha,-\alpha\}$-code, then we have already observed that $\left(M\odot\overline{M}\right)^{-1}=\frac{1}{1-\alpha^{2}}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)$ and so the inequality $\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{{\dagger}}\mathbbm{1}\leq r$ in the above is equivalent to the relative bound.
Remark 2.8.
If one is interested in efficiently computing the term $\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{{\dagger}}\mathbbm{1}$ in Theorem 2.6 when $M\odot\overline{M}$ is invertible, then this can be done without having to determine $\left(M\odot\overline{M}\right)^{{\dagger}}=\left(M\odot\overline{M}\right)^{-1}$. Indeed, it suffices to find a solution $x\in\mathbb{R}^{n}$ to $\left(M\odot\overline{M}\right)x=\mathbbm{1}$, since then $\mathbbm{1}^{\intercal}x=\mathbbm{1}^{\intercal}\left(M\odot\overline{M}\right)^{-1}\mathbbm{1}$.
Connection to Delsarte’s linear programming method
We conclude this section by discussing the connection between our projection based method and the classical linear programming approach of Delsarte, Goethels, and Seidel [28, 29] via Gegenbauer polynomials. Given a function $f\colon\mathbb{R}\rightarrow\mathbb{R}$ and a matrix $M\in\mathbb{R}^{n\times n}$, we let $f(M)\in\mathbb{R}^{n\times n}$ denote the matrix resulting from applying $f$ to each entry of $M$, i.e. $f(M)(i,j)=f(M(i,j))$. Observe that if we let $\mathscr{C}=\{v_{1},\ldots,v_{n}\}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$, then with respect to the Frobenius inner product, projecting each $v_{i}v_{i}^{\intercal}$ onto the orthogonal complement of the identity $I$ results in a collection of matrices $\mathscr{C}^{\prime}=\{v_{1}v_{1}^{\intercal}-\frac{1}{r}I,\ldots,v_{n}v_{n}^{\intercal}-\frac{1}{r}I\}$ which has the Gram matrix $M_{\mathscr{C}^{\prime}}=G^{r}_{2}(M_{\mathscr{C}})$, where $G^{r}_{2}(x)=x^{2}-\frac{1}{r}$ is a scaled version of the second Gegenbauer polynomial of the $(r-1)$-sphere.
While the Gegenbauer polynomials and their properties are usually derived using orthogonal spaces of harmonic homogeneous polynomials, see e.g. [29], one may equivalently obtain them via an appropriate generalization of the previous argument to symmetric tensors, see Ehrentraut and Muschik [33] for more information. Indeed, a symmetric matrix which is orthogonal to the identity can be viewed as a traceless symmetric 2-tensor and in general, there is an isomorphism between the inner product spaces of traceless symmetric $k$-tensors and harmonic homogeneous polynomials of degree $k$. While this approach to spherical harmonics via symmetric tensors does not seem to be common in the mathematics literature, it is more well known in the physics literature, see e.g. [75, 54, 33, 92, 23].
In addition to being generalizations of the Chebyshev and Legendre polynomials, the significance of Gegenbauer polynomials goes back to a classical result of Schoenberg [81], who proved that a function $f\colon[-1,1]\rightarrow\mathbb{R}$ has the property that $f(M_{\mathscr{C}})$ is positive semidefinite for any finite set of unit vectors $\mathscr{C}$ in $\mathbb{R}^{r}$ if and only if $f$ is a nonnegative linear combination of Gegenbauer polynomials. Indeed, the inequality $\mathbbm{1}^{\intercal}f(M_{\mathscr{C}})\mathbbm{1}\geq 0$ with $f$ chosen appropriately [^6] If $f$ is a nonnegative linear combination of Gegenbauer polynomials such that $f(\left\langle v_{i},v_{j}\right\rangle)\leq-1$ for all $i\neq j$, then $0\leq\mathbbm{1}^{\intercal}f(M_{\mathscr{C}})\mathbbm{1}\leq nf(1)-n(n-1)$, so that $n\leq f(1)+1$. Therefore, to get the best bound on $n$, we must choose $f$ subject to the given constraints so as to minimize $f(1)$. This minimization can be done using linear programming. underlies the linear programming approach of Delsarte, Goethels, and Seidel, which has the benefit of only requiring an upper bound on the off-diagonal entries of $f(M_{\mathscr{C}})$ in order to obtain an upper bound on $n$. Our approach could therefore be seen as a more refined use of the Frobenius inner product geometry of (symmetric) matrices/tensors in order to obtain sharp bounds on the largest eigenvalue of $M_{\mathscr{C}}$ which then yield sharp upper bounds on $n$. Although one could obtain the same results as we do by working with polynomials with respect to a certain inner product, we believe that using symmetric tensors is both simpler and more suggestive. The drawback of our approach seems to be that it require some control over the (generalized) inverse of $f(M_{\mathscr{C}})$ in order to be effective.
3 Equiangular lines in $\mathbb{R}^{r}$
In this section we will prove our main theorems regarding real equiangular lines. Our methods build on some of the ideas appearing in [7, 59, 61]. We will first give an outline for proving Theorem 1.2. To this end, we begin by establishing some basic definitions and observations.
Note that the case $\alpha=0$ is trivial since it corresponds to orthogonal vectors, so we can assume that $\alpha>0$. For any family of equiangular lines $\mathscr{L}$, we say that a spherical $\{\alpha,-\alpha\}$-code $\mathscr{C}$ represents $\mathscr{L}$ if there exists a bijection $\ell\colon\mathscr{C}\rightarrow\mathscr{L}$ such that $v\in\ell(v)$ for all $v\in\mathscr{C}$. As noted in previous sections, choosing a unit vector along each line $l\in\mathscr{L}$ always yields a spherical $\{\alpha,-\alpha\}$-code which represents $\mathscr{L}$, where $\alpha$ is the cosine of the common angle. Therefore, instead of working with a family of lines directly, we will always consider some spherical $\{\alpha,-\alpha\}$-code $\mathscr{C}$ which represents it.
Our first lemma shows that we can assume without loss of generality that an eigenvector corresponding to the largest eigenvalue of the corresponding Gram matrix $M_{\mathscr{C}}$ has no negative coordinates. It is proved by negating a suitable subset of vectors. Indeed, this operation is known in the literature as switching, and it is useful for us because it negates coordinates of eigenvectors of $M_{\mathscr{C}}$ without affecting eigenvalues.
Lemma 3.1.
For any family $\mathscr{L}$ of equiangular lines, there exists a spherical $\{\alpha,-\alpha\}$-code $\mathscr{C}$ which represents $\mathscr{L}$ such that an eigenvector corresponding to the largest eigenvalue of the Gram matrix $M_{\mathscr{C}}$ has no negative coordinates.
Proof.
Let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code which represents $\mathscr{L}$ and let $x$ be an eigenvector of $M_{\mathscr{C}}$ corresponding to $\lambda_{1}(M_{\mathscr{C}})$. Let $S=\{u:x(u)<0\}$ denote the negative coordinates of $x$ and let $D$ be the $\mathscr{C}\times\mathscr{C}$ diagonal matrix given by $D(u,u)=-1$ if $u\in S$ and $D(u,u)=1$ otherwise, so that $Dx$ has no negative coordinates. Now let $\mathscr{C}^{\prime}=(\mathscr{C}\backslash S)\cup\{-u:u\in S\}$ be the spherical $\{\alpha,-\alpha\}$-code obtained from $\mathscr{C}$ by negating the vectors corresponding to $S$, and note that it clearly also represents $\mathscr{L}$. Since its Gram matrix satisfies $M_{\mathscr{C}^{\prime}}=DM_{\mathscr{C}}D$, it is easy to check that $Dx$ is an eigenvector of $M_{\mathscr{C}^{\prime}}$ corresponding to $\lambda_{1}(M_{\mathscr{C}})=\lambda_{1}(M_{\mathscr{C}^{\prime}})$. ∎
As previously discussed, we can associate a corresponding graph $G_{\mathscr{C}}$ to any spherical $\{\alpha,-\alpha\}$-code $\mathscr{C}$ with an edge for each pair of vectors having a negative inner product. More precisely, we define $G=G_{\mathscr{C}}$ to have vertex set $V(G)=\mathscr{C}$ and edge set $E(G)=\{uv:u,v\in\mathscr{C}\text{ and }\left\langle u,v\right\rangle=-\alpha\}$. Moreover, letting $A=A_{G}$ be the adjacency matrix of $G$ and $M=M_{\mathscr{C}}$ be the Gram matrix of $\mathscr{C}$, we will make repeated use of the fact that
$M=(1-\alpha)I+\alpha J-2\alpha A.$ (2)
Before giving the outline of our approach for proving Theorem 1.2, we also observe that the upper bound of $\binom{1/\alpha^{2}-1}{2}$ on the size of a spherical $\{\alpha,-\alpha\}$-code follows immediately from the assumption that the largest eigenvalue of its Gram matrix is at most $\frac{1-\alpha^{2}}{2\alpha^{2}}$, and we also characterize when equality occurs.
Lemma 3.2.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code. If the largest eigenvalue $\lambda_{1}$ of the Gram matrix $M=M_{\mathscr{C}}$ satisfies $\lambda_{1}\leq\frac{1-\alpha^{2}}{2\alpha^{2}}$, then
$|\mathscr{C}|\leq\binom{1/\alpha^{2}-1}{2}$
with equality only if the span of $\mathscr{C}$ has dimension $21/\alpha^{2}-$ 2.
Proof.
Letting $n=|\mathscr{C}|$ and $r=\operatorname{rk}(M)$, the desired bound follows immediately from
$n(\alpha^{2}n+1-\alpha^{2})=\operatorname{tr}(M^{2})=\sum_{i=1}^{r}{\lambda_{i}(M)^{2}}\leq\lambda_{1}\operatorname{tr}(M)\leq\frac{1-\alpha^{2}}{2\alpha^{2}}n.$
Moreover, observe that equality in the above occurs only if $\lambda_{i}(M)=\frac{1-\alpha^{2}}{2\alpha^{2}}$ for all $1\leq i\leq r$. Note also that $r$ is the dimension of the span of $\mathscr{C}$ and by changing bases, we may assume that $\mathscr{C}\subset\mathbb{R}^{r}$, so that if we let $V\in\mathbb{R}^{r\times\mathscr{C}}$ be the matrix given by $Vy=\sum_{v\in\mathscr{C}}{y(v)v}$, then $M=V^{\intercal}V$ has the same nonzero eigenvalues as the $r\times r$ matrix $VV^{\intercal}=\sum_{v\in\mathscr{C}}{vv^{\intercal}}$. Therefore, $n=\binom{1/\alpha^{2}-1}{2}$ implies that $\sum_{v\in\mathscr{C}}{vv^{\intercal}}=\frac{1-\alpha^{2}}{2\alpha^{2}}I$ where $I$ is the $r\times r$ identity matrix and by taking the trace, we conclude that $\binom{1/\alpha^{2}-1}{2}=n=\frac{1-\alpha^{2}}{2\alpha^{2}}r$ and so $r=1/\alpha^{2}-2$. ∎
Outline of the proof of Theorem 1.2.
We may assume $\alpha\leq 1/7$ since otherwise the theorem follows from the results of Neumann, Lemmens and Seidel, and Cao, Koolen, Lin, and Yu [67, 19]. Now let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with $n=|\mathscr{C}|=N^{\mathbb{R}}_{\alpha}(r)$ as given by Lemma 3.1, so that the corresponding Gram matrix $M=M_{\mathscr{C}}$ has largest eigenvalue $\lambda_{1}$ with a corresponding unit eigenvector $x$ having no negative coordinates. Note that $(2+3\alpha)r$ is larger than $\binom{1/\alpha^{2}-1}{2}$ precisely when $r>r_{\alpha}=\binom{1/\alpha^{2}-1}{2}/(2+3\alpha)$ and so our goal is will be to bound $n$ from above by $\max\left(\binom{1/\alpha^{2}-1}{2},(2+3\alpha)r\right)$. As such, we may also assume without loss of generality that $n\geq\binom{1/\alpha^{2}-1}{2}$. Finally, if $\lambda_{1}\leq\frac{1-\alpha^{2}}{2\alpha^{2}}$ then we obtain the first desired upper bound of $\binom{1/\alpha^{2}-1}{2}$ using Lemma 3.2.
Otherwise, we will have $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$ in which case Theorem 2.1 will imply that $\lambda_{1}$ is actually much larger (nearly $\alpha n$). More precisely, we will show that $x(u)>\alpha\sqrt{n}/\left(\lambda_{1}+\frac{1-\alpha^{2}}{4\alpha^{2}}\right)$ for any $u\in\mathscr{C}$, so that together with (2), we will conclude [^7] Intuitively, our argument works because it is implicitly showing that $x$ must be close to the normalized all ones vector $\frac{1}{\sqrt{n}}\mathbbm{1}$. that the maximum degree $\Delta$ of the corresponding graph satisfies $\Delta<\frac{1+3\alpha^{2}}{8\alpha^{3}}$. We will then apply a bootstrapping argument [^8] It follows from (2) that $\lambda_{2}(G)\leq\frac{1-\alpha}{2\alpha}$. If we knew that $G$ was $\Delta$-regular and sufficiently sparse, then the usual Alon–Boppana theorem would yield the desired bound $\frac{1}{2\alpha}\geq\sqrt{\Delta}$. via a variant of the usual Alon–Boppana theorem in order to conclude that $\Delta\leq\frac{1}{4\alpha^{2}}$. Also via (2), it will follow that $\lambda_{1}\geq\frac{\mathbbm{1}^{\intercal}M\mathbbm{1}}{n}\geq 1-\alpha+\alpha n-\frac{1}{2\alpha}$, so that $\lambda_{1}$ will indeed be almost as large as $\alpha n$, which is equivalent to the matrix $L=M-\lambda_{1}xx^{\intercal}$ having a small Frobenius norm. On the other hand, $L$ has a large trace (roughly $n$), and so we will conclude that $L$ has large rank (roughly $n/2$). The second desired bound $n\leq(2+3\alpha)r$ will then follow since the rank of $L$ is at most $r-1$. ∎
Remark 3.3.
The proof of Theorem 1.4 will be almost identical to that of Theorem 1.2, except that we will apply the bootstrapping argument with a stronger version of the Alon-Boppana theorem, as done in [59]. Furthermore, Theorem 1.5 will follow by combining our new bound on the maximum degree with the second eigenvalue multiplicity argument of [61].
In order to implement the arguments in the outline given above, we now state some more useful definitions and results. Since we will need them in Section 4, we derive the more general complex versions of some of these results. We begin by giving a simple upper bound on the largest eigenvalue of the corresponding Gram matrix.
Lemma 3.4.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{C}^{r}$ with corresponding Gram matrix $M=M_{\mathscr{C}}$. Then $\lambda_{1}(M)\leq\alpha|\mathscr{C}|+1-\alpha$.
Proof.
Let $x$ be a unit eigenvector corresponding to $\lambda_{1}(M)$. Using Cauchy–Schwarz, we obtain
$\lambda_{1}(M)=x^{\intercal}Mx\leq\sum_{u,v\in\mathscr{C}}{|M(u,v)||x(u)||x(v)|}$ $=(1-\alpha)\sum_{v\in\mathscr{C}}{|x(v)|^{2}}+\alpha\left(\sum_{v\in\mathscr{C}}{|x(v)|}\right)^{2}$
$\leq(1-\alpha)\sum_{v\in\mathscr{C}}{|x(v)|^{2}}+\alpha|\mathscr{C}|\sum_{v\in\mathscr{C}}{|x(v)|^{2}}$
$=1-\alpha+\alpha|\mathscr{C}|.\qed$
We will also need the following lemma which shows that a matrix with small Frobenius norm and large trace must have high rank. Note that this result, whose proof is a simple application of Cauchy–Schwarz, was an important ingredient of our arguments in [7] and has many other applications, see e.g. the survey of Alon [2].
Lemma 3.5.
For any Hermitian matrix $L$, we have
$\operatorname{tr}(L)^{2}\leq\operatorname{tr}(L^{2})\operatorname{rk}(L).$
Proof.
Let $r=\operatorname{rk}(L)$ and note that $\operatorname{tr}(L)=\sum_{i=1}^{r}{\lambda_{i}(L)}$ and $\operatorname{tr}(L^{2})=\sum_{i=1}^{r}{\lambda_{i}(L)^{2}}$. The result now follows via Cauchy–Schwarz. ∎
Remark 3.6.
In keeping with the theme of this paper, we note that one can alternatively prove Lemma 3.5 using the Frobenius inner product, under the condition that $L$ is positive semidefinite (which will indeed be the case in our applications). Indeed, if $L$ is $n\times n$ with rank $r$, then there exists an $r\times n$ matrix $V$ such that $L=V^{*}V$ and so if we let $I$ be the $r\times r$ identity matrix, then applying Cauchy–Schwarz with the Frobenius inner product yields
$\operatorname{tr}(L)^{2}=\operatorname{tr}(VV^{*})^{2}=\left\langle VV^{*},I\right\rangle_{F}^{2}\leq||VV^{*}||_{F}^{2}||I||_{F}^{2}=\operatorname{tr}(VV^{*}VV^{*})\operatorname{tr}(I)=\operatorname{tr}(L^{2})r.$
We now show how Lemma 3.4 and Lemma 3.5 can be used to obtain an upper bound on the size of a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$, in terms of $r$, $\alpha$, and the average degree $\overline{d}$ of the corresponding graph.
Lemma 3.7.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ such that the corresponding graph $G=G_{\mathscr{C}}$ has average degree $\overline{d}=\overline{d_{G}}$. Then
$|\mathscr{C}|<\left(1+\left(\frac{2\alpha}{1-\alpha}\right)^{2}\overline{d}\right)(r-1)+2.$
Proof.
Let $n=|\mathscr{C}|$ and let $M=M_{\mathscr{C}}$ be the corresponding Gram matrix having largest eigenvalue $\lambda_{1}$ with $x$ being a corresponding unit eigenvector. We have via (2) that $\lambda_{1}\geq\frac{\mathbbm{1}^{\intercal}M\mathbbm{1}}{n}=1-\alpha+\alpha n-2\alpha\overline{d}$ and thus
$\lambda_{1}^{2}\geq\alpha^{2}n^{2}-\left(4\alpha^{2}\overline{d}-2\alpha(1-\alpha)\right)n+\left(2\alpha\overline{d}-1+\alpha\right)^{2}\geq\alpha^{2}n^{2}-\left(4\alpha^{2}\overline{d}-2\alpha+2\alpha^{2}\right)n.$
Therefore, the matrix $L=M-\lambda_{1}xx^{\intercal}$ satisfies
$\operatorname{tr}(L^{2})=||M||_{F}^{2}-\lambda_{1}^{2}=\alpha^{2}n^{2}+(1-\alpha^{2})n-\lambda_{1}\leq\left((1-\alpha)^{2}+4\alpha^{2}\overline{d}\right)n.$
On the other hand, Lemma 3.4 implies $\operatorname{tr}(L)=n-\lambda_{1}\geq(1-\alpha)(n-1)$ and moreover we have $\operatorname{rk}(L)=r-1$. Therefore, applying Lemma 3.5 yields the desired bound
$n-2<\frac{(n-1)^{2}}{n}\leq\frac{\operatorname{tr}(L)^{2}}{(1-\alpha)^{2}n}\leq\frac{\operatorname{tr}(L^{2})}{(1-\alpha)^{2}n}\operatorname{rk}(L)\leq\left(1+\frac{4\alpha^{2}}{(1-\alpha)^{2}}\overline{d}\right)(r-1).\qed$
In view of Lemma 3.7, we now turn our attention toward obtaining bounds on the maximum degree $\Delta$ of the corresponding graph. We first apply Theorem 2.1 to obtain an upper bound on $\Delta$, provided that $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$ and $n$ is sufficiently large.
Lemma 3.8.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code with $|\mathscr{C}|\geq\binom{1/\alpha^{2}-1}{2}$. If the largest eigenvalue $\lambda_{1}=\lambda_{1}(M)$ of the Gram matrix $M=M_{\mathscr{C}}$ satisfies $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$ and has a corresponding eigenvector $x$ with no negative coordinates, then the maximum degree of the corresponding graph $G=G_{\mathscr{C}}$ satisfies
$\Delta_{G}<\frac{1+3\alpha^{2}-4\alpha^{3}}{8\alpha^{3}}.$
Proof.
By normalizing, we may assume that $x$ is a unit vector. Now let $n=|\mathscr{C}|$, let $u\in\mathscr{C}$, and observe that Theorem 2.1 implies
$\frac{x(u)^{2}}{n}\geq\frac{1-\frac{1-\alpha^{2}}{2\alpha^{2}\lambda_{1}}}{\frac{n\lambda_{1}^{2}}{\alpha^{2}n+1-\alpha^{2}}-\frac{1-\alpha^{2}}{2\alpha^{2}}n}\eqqcolon Q$
where the numerator and denominator of $Q$ must be positive. Since $x$ has no negative coordinates, we conclude that $x(u)\geq\sqrt{Qn}$ for all $u\in\mathscr{C}$ and it follows from (2) that
$(\lambda_{1}-1+\alpha)\sqrt{Qn}\leq(\lambda_{1}-1+\alpha)x(u)=((M-(1-\alpha)I)x)(u)$ $=((\alpha J-2\alpha A)x)(u)$
$=\alpha\left\langle\mathbbm{1},x\right\rangle-2\alpha\sum_{v\in N_{G}(u)}{x(v)}$
$\leq\alpha\sqrt{n}-2\alpha\sqrt{Qn}\cdot d_{G}(u),$
so we have $d_{G}(u)\leq\frac{1}{2\sqrt{Q}}-\frac{\lambda_{1}}{2\alpha}+\frac{1-\alpha}{2\alpha}$. It now remains to verify that $Q$ satisfies the lower bound
$Q>\frac{\alpha^{2}}{\lambda_{1}^{2}\left(1+\frac{1-\alpha^{2}}{2\alpha^{2}\lambda_{1}}\right)}.$ (3)
Indeed, if (3) holds then we have $\frac{1}{\sqrt{Q}}<\frac{\lambda_{1}}{\alpha}\sqrt{1+\frac{1-\alpha^{2}}{2\alpha^{2}\lambda_{1}}}<\frac{\lambda_{1}}{\alpha}+\frac{1-\alpha^{2}}{4\alpha^{3}}$ and thus
$d_{G}(u)<\frac{1-\alpha^{2}}{8\alpha^{3}}+\frac{1-\alpha}{2\alpha}=\frac{1+3\alpha^{2}-4\alpha^{3}}{8\alpha^{3}}.$
Now observe that by cross multiplying and collecting terms, (3) is equivalent to
$\left(\frac{1-\alpha^{2}}{2\alpha^{2}}\right)^{2}-\frac{1-\alpha^{2}}{2}n<\lambda_{1}^{2}\left(1-\frac{\alpha^{2}n}{\alpha^{2}n+1-\alpha^{2}}\right)=\frac{1-\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}\lambda_{1}^{2},$
so if we define $f(n)=\left(\frac{1-\alpha^{2}}{4\alpha^{4}}-\frac{n}{2}\right)\left(\alpha^{2}n+1-\alpha^{2}\right)$, then this is also equivalent to $f(n)<\lambda_{1}^{2}$. By computing its derivative, it is easy to see that $f(n)$ is decreasing for $n\geq\frac{1}{2}\binom{1/\alpha^{2}-1}{2}$ and we have $n\geq\max\left(\binom{1/\alpha^{2}-1}{2},0\right)\geq\frac{1}{2}\binom{1/\alpha^{2}-1}{2}$, so $f$ is maximized when $n=\max\left(\binom{1/\alpha^{2}-1}{2},0\right)$. Now (3) follows by observing that for both $n=\binom{1/\alpha^{2}-1}{2}$ and $n=0$, we have $f(n)=\left(\frac{1-\alpha^{2}}{2\alpha^{2}}\right)^{2}<\lambda_{1}^{2}$. ∎
Remark 3.9.
Lemma 3.8 together with Lemma 3.7 and Lemma 3.2 already imply that $N^{\mathbb{R}}_{\alpha}(r)\leq\max\left(\binom{1/\alpha^{2}-1}{2},O\left(r/\alpha\right)\right)$. Moreover, instead of using Lemma 3.8, we note that a more precise application of Theorem 2.1 can be used to show that the $O\left(r/\alpha\right)$ term is at most $\frac{1+\alpha}{2\alpha}r+O(1/\alpha^{3})$ (see Section 4 for the analogous argument in the complex case). Such a bound was recently conjectured to be obtainable from the second level of the Lasserre hierarchy by de Laat, Keizer, and Machado [66], so it would be interesting to determine if this conjecture follows from our arguments.
To improve on the $O(r/\alpha)$ bound mentioned in the preceding remark, we will strengthen the degree bound of Lemma 3.8 by a factor of $\alpha$ by using a bootstrapping argument via an Alon–Boppana-type theorem. To this end, it will be convenient to introduce the following spectral graph parameter. For any graph $G$ with adjacency matrix $A=A_{G}$, we define
$\beta(G)=\max_{\begin{subarray}{c}x\in\mathbb{R}^{V(G)}\backslash\{0\}\\ x\perp\mathbbm{1}\end{subarray}}{\frac{x^{\intercal}Ax}{x^{\intercal}x}},$
i.e. the maximum Rayleigh quotient over all vectors orthogonal to the all ones vector $\mathbbm{1}$. Note that via the Courant min-max principle, the second largest eigenvalue satisfies $\lambda_{2}(G)\leq\beta(G)$ with equality when $G$ is regular, so that $\beta$ can be seen as an extension to all graphs of the second largest eigenvalue of a regular graph. Using (2), we immediately obtain the following characterization of $\beta$ for the graph corresponding to a spherical $\{\alpha,-\alpha\}$-code.
Lemma 3.10.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with corresponding graph $G=G_{\mathscr{C}}$. Then
$\beta(G)\leq\frac{1-\alpha}{2\alpha}$
with equality whenever $|\mathscr{C}|\geq r+2$.
Proof.
Let $M$ and $A$ be the corresponding Gram and adjacency matrices, respectively and let $x\in\mathbb{R}^{\mathscr{C}}\backslash\{0\}$ be such that $x\perp\mathbbm{1}$. Then $Jx=0$ and so via (2) we have
$0\leq x^{\intercal}Mx=(1-\alpha)x^{\intercal}Ix-2\alpha x^{\intercal}Ax,$
so that $\frac{x^{\intercal}Ax}{x^{\intercal}x}\leq\frac{1-\alpha}{2\alpha}$. Since $x$ was arbitrary, we conclude that $\beta(G)\leq\frac{1-\alpha}{2\alpha}$ as desired.
Moreover, if $|\mathscr{C}|\geq r+2$ then since $\operatorname{rk}(M)\leq r$, the nullspace of $M$ has dimension at least 2 and must therefore contain a nonzero vector $y$ which is orthogonal to $\mathbbm{1}$. Using (2), we conclude
$0=y^{\intercal}My=(1-\alpha)y^{\intercal}y-2\alpha y^{\intercal}Ay,$
which implies that $\beta(G)\geq\frac{y^{\intercal}Ay}{y^{\intercal}y}=\frac{1-\alpha}{2\alpha}$. ∎
Having established an upper bound on $\beta$, we now state a variant of the Alon–Boppana theorem which gives a lower bound on $\beta$ for any graph $G$ with a given maximum degree. In particular, we follow Friedman and Tillich’s approach [40] of starting with an eigenvector corresponding to a subgraph $H\subseteq G$ and then “projecting out the constant” in order to make this eigenvector orthogonal to the all ones vector.
Lemma 3.11.
Let $G$ be a graph on $n$ vertices with maximum degree $\Delta=\Delta(G)$. For any subgraph $H\subseteq G$ with $x$ being a unit eigenvector corresponding to $\lambda_{1}(H)$, we have
$\beta(G)\geq\lambda_{1}(H)-\frac{2\Delta}{n}\left\langle x,\mathbbm{1}\right\rangle^{2}.$
Proof.
First note that by the Perron–Frobenius theorem, all coordinates of $x$ are nonnegative. Now observe that we may extend $x$ to $\mathbb{R}^{V(G)}$ by defining $x(u)=0$ for all $u\notin V(H)$, so that we conclude $x^{\intercal}Ax\geq\lambda_{1}(H)$. Now we define $y=x-\frac{\left\langle x,\mathbbm{1}\right\rangle}{n}\mathbbm{1}$, i.e. the projection of $x$ onto the orthogonal complement of $\mathbbm{1}$. The desired bound now follows by computing that $y^{\intercal}y=||x||^{2}-\frac{\left\langle x,\mathbbm{1}\right\rangle^{2}}{n}\leq 1$ and
$y^{\intercal}Ay=x^{\intercal}Ax-2\frac{\left\langle x,\mathbbm{1}\right\rangle}{n}\mathbbm{1}^{\intercal}Ax+\frac{\left\langle x,\mathbbm{1}\right\rangle^{2}}{n^{2}}\mathbbm{1}^{\intercal}A\mathbbm{1}$ $\geq\lambda_{1}(H)-2\frac{\left\langle x,\mathbbm{1}\right\rangle}{n}\sum_{u\in V(G)}{x(u)d_{G}(u)}$
$\geq\lambda_{1}(H)-\frac{2\Delta}{n}\left\langle x,\mathbbm{1}\right\rangle^{2}.\qed$
The previous lemma suggests finding a small subgraph $H$ with $\lambda_{1}(H)$ large and indeed, in the proof of the usual Alon–Boppana theorem one takes $H$ to be a ball centered at a vertex. Indeed, for the proof of Theorem 1.2 we take $H=K_{1,t}$, i.e. the graph consisting of a vertex connected to $t$ other vertices, and so we will first determine $\lambda_{1}(K_{1,t})$ and its corresponding unit eigenvector.
Lemma 3.12.
For all $t\in\mathbb{N}$, we have $\lambda_{1}(K_{1,t})=\sqrt{t}$ and the corresponding unit eigenvector $x$ satisfies $\left\langle\mathbbm{1},x\right\rangle=(\sqrt{t}+1)/\sqrt{2}$.
Proof.
Let $H=K_{1,t}$ so that we have $V(H)=T\cup\{w\}$ and $E(H)=\{wv:v\in T\}$. If we let $A=A_{H}$ be the corresponding adjacency matrix, then it is a straightforward calculation to verify that the maximum of $x^{\intercal}Ax=2x(w)\sum_{v\in T}{x(v)}$ over all unit vectors $x\in\mathbb{R}^{V(H)}$ occurs when $x(v)=1/\sqrt{2t}$ for $v\in T$ and $x(w)=1/\sqrt{2}$, in which case we have $x^{\intercal}Ax=\sqrt{t}$ and $\left\langle\mathbbm{1},x\right\rangle=(\sqrt{t}+1)/\sqrt{2}$. ∎
We now combine Lemma 3.11 and Lemma 3.12 to obtain the desired lower bound on $\beta$.
Lemma 3.13.
Let $G$ be a graph on $n$ vertices with maximum degree $\Delta=\Delta_{G}$. For all $t\in\mathbb{N}$ with $t\leq\Delta$, we have
$\beta(G)\geq\sqrt{t}-\frac{\Delta}{n}\left(\sqrt{t}+1\right)^{2}.$
Proof.
Let $w\in V(G)$ be a vertex having degree $d_{G}(v)=\Delta$ and let $H$ be a subgraph consisting of $w$ together with $t$ of its neighbors. Applying Lemma 3.12 and Lemma 3.11 with $H$, we have
$\beta(G)\geq\lambda_{1}(H)-\frac{2\Delta}{n}\frac{\left(\sqrt{t}+1\right)^{2}}{2}\geq\sqrt{t}-\frac{\Delta}{n}\left(\sqrt{t}+1\right)^{2}.\qed$
Using Lemma 3.10 and Lemma 3.13, we now derive the bootstrapping argument which improves the upper bound on the maximum degree of Lemma 3.8 by a factor of $\alpha$.
Lemma 3.14.
Let $70<\alpha\leq 1/$ 7 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code with $|\mathscr{C}|\geq\binom{1/\alpha^{2}-1}{2}$. If the largest eigenvalue $\lambda_{1}$ of the Gram matrix $M=M_{\mathscr{C}}$ satisfies $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$ and has a corresponding eigenvector $x$ with no negative coordinates, then the maximum degree of the corresponding graph $G=G_{\mathscr{C}}$ satisfies
$\Delta_{G}<\frac{1}{4\alpha^{2}}.$
Proof.
Let $n=|\mathscr{C}|$. Using Lemma 3.8, we have $\Delta=\Delta_{G}<\frac{1+3\alpha^{2}}{8\alpha^{3}}$. Since $n\geq\binom{1/\alpha^{2}-1}{2}>\frac{1-3\alpha^{2}}{2\alpha^{4}}$ and $\alpha\leq 1/7$, we conclude that
$\frac{\Delta}{n}\leq\frac{(1+3\alpha^{2})\alpha}{(1-3\alpha^{2})4}<\frac{\alpha}{3}<\frac{1}{20}.$
We will now show that $\Delta\leq\left\lceil\left(\frac{n}{4\Delta}\right)^{2}\right\rceil-1$, so suppose for sake of contradiction that $\Delta\geq\left\lceil\left(\frac{n}{4\Delta}\right)^{2}\right\rceil$. Let $s=\left\lceil\left(\frac{n}{4\Delta}\right)^{2}\right\rceil$ and observe that $\sqrt{s}<\sqrt{\left(\frac{n}{4\Delta}\right)^{2}+1}<\frac{n}{4\Delta}+\frac{1}{10}$, so that applying Lemma 3.10 and Lemma 3.13 with $t=s$ yields
$\frac{1-\alpha}{2\alpha}\geq\beta(G)\geq\sqrt{s}-\frac{\Delta}{n}\left(\sqrt{s}+1\right)^{2}\geq\frac{n}{4\Delta}-\frac{\Delta}{n}\left(\frac{n^{2}}{16\Delta^{2}}+\frac{11n}{20\Delta}+\frac{121}{100}\right)$ $=\frac{3n}{16\Delta}-\frac{11}{20}-\frac{121\Delta}{100n}$
$>\frac{9}{16\alpha}-\frac{11}{20}-\frac{121\alpha}{300},$
which contradicts the assumption that $\alpha\leq 1/7$.
Now that we have established $\Delta\leq\left\lceil\left(\frac{n}{4\Delta}\right)^{2}\right\rceil-1<\left(\frac{n}{4\Delta}\right)^{2}$, we have $\sqrt{\Delta}<\frac{n^{1/3}}{2^{2/3}}$ and so we apply Lemma 3.10 and Lemma 3.13 once again, but this time with $t=\Delta$, in order to conclude that
$\frac{1-\alpha}{2\alpha}\geq\beta(G)\geq\sqrt{\Delta}\left(1-\frac{\sqrt{\Delta}}{n}\left(\sqrt{\Delta}+1\right)^{2}\right)$ $\geq\sqrt{\Delta}\left(1-\frac{n^{1/3}}{2^{2/3}n}\left(\frac{n^{1/3}}{2^{2/3}}+1\right)^{2}\right)$
$=\sqrt{\Delta}\left(1-\frac{1}{4}-\frac{1}{(2n)^{1/3}}-\frac{1}{(2n)^{2/3}}\right)$
$>\frac{2}{3}\sqrt{\Delta},$
where the last inequality follows since $\alpha\leq 1/7$ implies that $n\geq\binom{1/\alpha^{2}-1}{2}\geq 1128$. Thus we have established that $\sqrt{\Delta}<\frac{3(1-\alpha)}{4\alpha}$ and again using $\alpha\leq 1/7$, we therefore compute
$\frac{\sqrt{\Delta}}{n}<\frac{3(1-\alpha)}{4\alpha}\frac{2\alpha^{4}}{(1-\alpha^{2})(1-2\alpha^{2})}=\frac{3\alpha^{3}}{2(1+\alpha)(1-2\alpha^{2})}<\frac{3\alpha^{3}}{2}$
and
$\Delta+2\sqrt{\Delta}+1<\frac{9(1-\alpha)^{2}}{16\alpha^{2}}+\frac{3(1-\alpha)}{2\alpha}+1=\frac{9}{16\alpha^{2}}+\frac{3}{8\alpha}+\frac{1}{16}<\frac{2}{3\alpha^{2}},$
The desired bound $\Delta<\frac{1}{4\alpha^{2}}$ now follows by applying Lemma 3.10 and Lemma 3.13 one final time with $t=\Delta$, yielding
$\frac{1-\alpha}{2\alpha}\geq\beta(G)\geq\sqrt{\Delta}\left(1-\frac{\sqrt{\Delta}}{n}\left(\Delta+2\sqrt{\Delta}+1\right)\right)>\sqrt{\Delta}(1-\alpha).\qed$
We now have all of the necessary ingredients to establish Theorem 1.2
Proof of Theorem 1.2.
We first consider the case $\alpha>1/7$. If $\alpha\notin\{1/3,1/5\}$, then Theorem 1.2 follows from Neumann’s bound $N^{\mathbb{R}}_{\alpha}(r)\leq 2r$ (see [67]) . Also if $r\leq 1/\alpha^{2}-2$ then $N^{\mathbb{R}}_{\alpha}(r)\leq\binom{1/\alpha^{2}-1}{2}$ follows from the relative bound. Otherwise if $\alpha\in\{1/3,1/5\}$ and $r\geq 1/\alpha^{2}-2$ then Theorem 1.2 follows from the theorems of Lemmens and Seidel [67] and Cao, Koolen, Lin, and Yu [19], which were mentioned in Section 1.1. Therefore, we henceforth assume that $\alpha\leq 1/7$.
As explained in the outline, we let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with $n=|\mathscr{C}|=N^{\mathbb{R}}_{\alpha}(r)$ given by Lemma 3.1, so that the Gram matrix $M=M_{\mathscr{C}}$ has largest eigenvalue $\lambda_{1}$ with a corresponding eigenvector $x$ having no negative coordinates. Since $(2+3\alpha)r$ is larger than $\binom{1/\alpha^{2}-1}{2}$ precisely when $r>r_{\alpha}=\binom{1/\alpha^{2}-1}{2}/(2+3\alpha)$, it will suffice to show that if $n\geq\binom{1/\alpha^{2}-1}{2}$, then exactly one of the following holds: Either $n\leq\binom{1/\alpha^{2}-1}{2}$ with equality only if $\mathscr{C}\subseteq\mathbb{R}^{1/\alpha^{2}-2}$ or otherwise $n\leq(2+3\alpha)r$. Indeed, these two cases will correspond to how large $\lambda_{1}$ is.
If $\lambda_{1}\leq\frac{1-\alpha^{2}}{2\alpha^{2}}$, then Lemma 3.2 yields the first desired bound $n\leq\binom{1/\alpha^{2}-1}{2}$, as well as a characterization of equality. Otherwise, $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$ and so we may apply Lemma 3.14 to conclude that the maximum degree of the corresponding graph $G=G_{\mathscr{C}}$ satisfies $\Delta_{G}<\frac{1}{4\alpha^{2}}$. Therefore, we may apply Lemma 3.7 to obtain the second desired bound
$n\leq\left(1+\frac{1}{(1-\alpha)^{2}}\right)(r-1)+2\leq\left(2+3\alpha\right)r,$
where the last inequality holds since $\alpha\leq 1/7$. ∎
In the previous proof, we used Lemma 3.13 with $H$ being a ball of radius 1, but in order to prove Theorem 1.4, we will take $H$ to be a ball of some radius $s\geq 2$. More specifically, for any graph $G$ with vertex $v\in V(G)$, we define $G(v,s)$ to be the subset of vertices for which there exists a path from $v$ of length at most $s$, i.e. the ball of radius $s$ centered at $v$ in $G$. We first recall the well-known fact that $|G(v,q)|\leq\sum_{i=0}^{s}{\Delta^{i}}=\frac{\Delta^{s+1}-1}{\Delta-1}$ where $\Delta=\Delta_{G}$ is the maximum degree. Indeed, this follows because there can be at most $\Delta^{i}$ vertices at a distance of exactly $i$ from $v$. We will also make use of the following lemma of Jiang [58] (see also Jiang and Polyanskii [59]) showing that for any $s\in\mathbb{N}$, every graph has a ball of radius $s$ with a large eigenvalue.
Lemma 3.15 (Jiang [58]).
Let $s\in\mathbb{N}$ and let $G$ be a graph on $n$ vertices with average degree $\overline{d}=\overline{d}(G)\geq 1$. Then there exists a vertex $v_{0}\in V(G)$ such that
$\lambda_{1}(G(v_{0},s))\geq 2\sqrt{\overline{d}-1}\cos\left(\frac{\pi}{s+2}\right).$
Proof of Theorem 1.4.
Via Lemma 3.1, let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with $n=|\mathscr{C}|=N^{\mathbb{R}}_{\alpha}(r)$ such that the Gram matrix $M=M_{\mathscr{C}}$ has a largest eigenvalue $\lambda_{1}$ with an eigenvector $x$ having no negative coordinates. As mentioned at the end of Section 1.1), by rotating each standard basis vector in $\mathbb{R}^{r}$ by the same angle towards the all ones vector $\mathbbm{1}$, it is easy to see that we must have $n\geq r$. Moreover, by assumption we have $r\gg 1/\alpha^{2s+1}\geq 1/\alpha^{5}$, so that Lemma 3.2 implies that $\lambda_{1}>\frac{1-\alpha^{2}}{2\alpha^{2}}$. It now follows from our previously established bounds that the average degree $\overline{d}$ and the maximum degree $\Delta$ of the corresponding graph satisfy $\overline{d}\leq\Delta\leq O(1/\alpha^{2})$. Indeed, if $\alpha>1/7$, then this follows from Lemma 3.8 and if $\alpha\leq 1/7$, then it follows from Lemma 3.14.
We now use Lemma 3.11 in order to obtain a stronger bound on $\overline{d}$. If $\overline{d}\leq 1$ then we are done, and otherwise we may apply Lemma 3.15 to obtain a vertex $v_{0}$ such that $H=G(v_{0},s)$, i.e. the ball of radius $s$ centered at $v_{0}$, satisfies $\lambda_{1}(H)\geq 2\sqrt{\overline{d}-1}\cos\left(\frac{\pi}{s+2}\right)$. Also, if we let $y$ be a unit eigenvector corresponding to $\lambda_{1}(H)$, then by Cauchy–Schwarz we have $\left\langle y,\mathbbm{1}\right\rangle^{2}\leq||y||^{2}\cdot||\mathbbm{1}||^{2}=|H|$ and since $|H|\leq\frac{\Delta^{s+1}-1}{\Delta-1}\leq O(\Delta^{s})$, we have that
$\frac{\Delta\left\langle y,\mathbbm{1}\right\rangle^{2}}{n}\leq O\left(\frac{\Delta^{s+1}}{r}\right)\leq O\left(\frac{1}{\alpha^{2s+2}r}\right)\leq\frac{o(1)}{\alpha}.$
Therefore, we may apply Lemma 3.10 and Lemma 3.11 in order to conclude
$\frac{1-\alpha}{2\alpha}\geq\beta(G)\geq\lambda_{1}(H)-\frac{2\Delta\left\langle y,\mathbbm{1}\right\rangle^{2}}{n}\geq 2\sqrt{\overline{d}-1}\cos\left(\frac{\pi}{s+2}\right)-\frac{o(1)}{\alpha},$
so that $\overline{d}\leq\left(\frac{1-\alpha}{2\alpha}\right)^{2}\cdot\frac{1+o(1)}{4\cos^{2}\left(\frac{\pi}{s+2}\right)}+1$. The desired result now follows via Lemma 3.7. ∎
We now turn to proving Theorem 1.5, for which we will need the following sublinear bound on the multiplicity of the second eigenvalue of a connected graph obtained by Jiang, Tidor, Yao, Zhang, and Zhao [61].
Theorem 3.16.
There exists a constant $B>0$ such that if $G$ is a connected graph on $n$ vertices with maximum degree $\Delta=\Delta(G)$, then the multiplicity of the second eigenvalue $\lambda_{2}=\lambda_{2}(G)$ satisfies
$\operatorname{m}_{G}(\lambda_{2})\leq\frac{B\log{\Delta}}{\log{\log{n}}}n.$
Proof.
In the proof of Theorem 2.2 in [61], one may take $c(\Delta,2)=\frac{1}{B\log{\Delta}}$ for a sufficiently large constant $B$. ∎
For the lower bound, we will need the following lemma. However, since its proof, as well as that of Theorem 1.5 follow directly from that of Proposition 3.2 and Theorem 1.2 of [61], respectively, we only provide sketches of the proofs here. In particular, we do not attempt to optimize the constant $C$ in Theorem 1.5.
Lemma 3.17.
Let $10<\alpha<$ 1 be such that the spectral radius order $k=k\left(\frac{1-\alpha}{2\alpha}\right)<\infty$. Then for all $r\in\mathbb{N}$, we have
$N^{\mathbb{R}}_{\alpha}(r)\geq\left\lfloor\frac{k}{k-1}(r-1)\right\rfloor.$
Proof sketch of Lemma 3.17.
By the definition of spectral radius order, there exists a graph $H$ with $k$ vertices such that $\lambda_{1}(H)=\frac{1-\alpha}{2\alpha}$, so we may let $G$ consist of $\left\lfloor\frac{r-1}{k-1}\right\rfloor$ disjoint copies of $H$ together with $d-1-(k-1)\left\lfloor\frac{r-1}{k-1}\right\rfloor$ isolated vertices. Note that $G$ has $n=\left\lfloor\frac{k(r-1)}{k-1}\right\rfloor$ vertices. Using the adjacency matrix of $A=A(G)$ of $G$, one may define $M=(1-\alpha)I+\alpha J-2\alpha A$ and observe that it is an $n\times n$ positive semidefinite matrix with rank at most $r$, so that it is the Gram matrix of some spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with $n$ elements. ∎
Proof sketch of Theorem 1.5.
First note that the lower bound in item 1 of Theorem 1.5 follows from Lemma 3.17. Now for the upper bounds, we begin as in previous proofs by using Lemma 3.1 to obtain a $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with $n=|\mathscr{C}|=N^{\mathbb{R}}_{\alpha}(r)$ such that the Gram matrix $M=M_{\mathscr{C}}$ has a largest eigenvalue with an eigenvector $x$ having no negative coordinates. We choose $C$ sufficiently large so that $r\geq 2^{1/\alpha^{C}}>\binom{1/\alpha^{2}-1}{2}$, and it is easy to see that $n\geq r$, so we may apply Lemma 3.2, and Lemma 3.8 to conclude that the corresponding graph has maximum degree $\Delta<1/\alpha^{3}$.
Since the nullspace of the corresponding Gram matrix $M=M_{\mathscr{C}}$ has dimension at least 2, using (2) one may conclude that $\frac{1-\alpha}{2\alpha}$ is an eigenvalue of $G$. Therefore $\lambda_{1}(G)\geq\frac{1-\alpha}{2\alpha}$ and so if we consider the connected components $G_{1},\ldots,G_{t}$ of $G$ ordered so that $\lambda_{1}(G_{1})\geq\ldots\geq\lambda_{1}(G_{t})$, then we either have $\lambda_{1}(G_{1})>\frac{1-\alpha}{2\alpha}$ or $\lambda_{1}(G_{1})=\frac{1-\alpha}{2\alpha}$.
If $\lambda_{1}(G_{1})>\frac{1-\alpha}{2\alpha}$, then one can show [^9] Using Lemma 3.10, this follows from $\frac{1-\alpha}{2\alpha}\geq\beta(G)\geq z^{\intercal}Az$ where $z$ is taken to be a linear combination of eigenvectors corresponding to $\lambda_{1}(G_{1})$ and $\lambda_{1}(G_{i})$, such that $z\perp\mathbbm{1}$ that $\lambda_{1}(G_{i})<\frac{1-\alpha}{2\alpha}$ for all $i\geq 2$. Therefore, $\operatorname{m}_{G_{1}}\left(\frac{1-\alpha}{2\alpha}\right)=\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)$. Moreover, using (2), the rank–nullity theorem, and the subadditivity of rank, one may show that $n-\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)\leq r+1$ and so we may apply Theorem 3.16 to $G_{1}$ to conclude
$n-r-1\leq\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)=\operatorname{m}_{G_{1}}\left(\frac{1-\alpha}{2\alpha}\right)\leq O\left(\frac{\log{\Delta}}{\log{\log{n}}}n\right)\leq O\left(\frac{\log(1/\alpha)}{\log{\log{r}}}n\right).$
Since $r\geq 2^{1/\alpha^{C}}$, we obtain the bound $n\leq r+\frac{C\log(1/\alpha)}{2\log{\log{r}}}r$ for $C$ sufficiently large.
Otherwise $\lambda_{1}(G_{1})=\frac{1-\alpha}{2\alpha}$, in which case the existence of $G_{1}$ implies that the spectral radius $k=k\left(\frac{1-\alpha}{2\alpha}\right)<\infty$, and so we have proven claim 3 of Theorem 1.5. Furthermore, one can show that $\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)\geq n-r+1$ (Indeed, by the rank–nullity theorem, it suffices to verify [^10] This follows from (2) by using the fact that $(1-\alpha)I-2\alpha A$ and $\alpha J$ are both positive semidefinite (so that the intersection of their nullspaces is the nullspace of $M$), together with the Perron–Frobenius theorem (there exists an eigenvector of $A$ corresponding to $\frac{1-\alpha}{2\alpha}$ with all nonnegative coordinates, so that it is not in the nullspace of $\alpha J$). that $\operatorname{rk}\left((1-\alpha)I-2\alpha A\right)\leq\operatorname{rk}(M)-1$). Now, if we let $S=\left\{i:\lambda_{1}(G_{i})=\frac{1-\alpha}{2\alpha}\right\}$, then by definition of $k$ we have $|V(G_{i})|\geq k$ for all $i\in S$ and hence $n\geq|S|k$. Moreover, the Perron–Frobenius theorem implies that $\lambda_{2}(G_{i})<\lambda_{1}(G_{i})$ for all $i$ and hence $|S|=\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)$. Therefore, $n\geq\operatorname{m}_{G}\left(\frac{1-\alpha}{2\alpha}\right)k\geq(n-r+1)k$ and so we conclude that $n\leq\left\lfloor\frac{k(r-1)}{k-1}\right\rfloor$.
Finally, observe that $r\geq 2^{1/\alpha^{C(k-1)}}$ implies $\left\lfloor\frac{k(r-1)}{k-1}\right\rfloor>r+\frac{C\log(1/\alpha)}{2\log{\log{r}}}r$ for $C$ sufficiently large, which establishes claim 1 of Theorem 1.5. Otherwise if $r<2^{1/\alpha^{C(k-1)}}$, then $\left\lfloor\frac{k(r-1)}{k-1}\right\rfloor<r+\frac{C\log(1/\alpha)}{\log{\log{r}}}r$ and so claim 2 of Theorem 1.5 follows. ∎
4 Equiangular lines in $\mathbb{C}^{r}$
The goal of this section will be to prove Theorem 1.11. The argument we will present is a generalization of part of the proof of Theorem 1.2, with Theorem 2.1 replaced by Theorem 2.4. The key difference will be that we do not use a bootstrapping [^11] Although one can generalize the definition of degree to the complex setting, we don’t know of an appropriate Alon–Boppana-type theorem which would work for this definition. argument and instead directly use Theorem 2.4 in order to obtain bounds on the largest eigenvalue of the corresponding Gram matrix, see the outline given below.
Now let $r\in\mathbb{N}$ and let $\mathscr{L}$ be a set of $n$ equiangular lines in $\mathbb{C}^{r}$. As previously noted, if we choose a complex unit vector along each line $\ell\in\mathscr{L}$, the resulting collection $\mathscr{C}$ forms a spherical $S^{1}(\alpha)$-code, where $S^{1}(\alpha)$ is the circle of radius $\alpha$ centered at 00 in $\mathbb{C}$. Similar to Section 3, we say that $\mathscr{C}$ represents $\mathscr{L}$ and so instead of working with $\mathscr{L}$ directly, we will consider some spherical $S^{1}(\alpha)$-code $\mathscr{C}$ which represents $\mathscr{L}$. Just as in Section 3, the case $\alpha=0$ is trivial and so we assume $\alpha>0$.
Outline of the proof of Theorem 1.11.
We will begin by summing the inequalities given by Theorem 2.4 over all $u\in\mathscr{C}$ in order to show that the largest eigenvalue $\lambda_{1}=\lambda_{1}(M)$ of the corresponding Gram matrix $M=M_{\mathscr{C}}$ satisfies $p(\lambda_{1})\geq 0$ where $p$ is a degree 3 polynomial with a positive leading term and 3 real roots $s_{0}\geq s_{1}>s_{2}$. It will follow that $\lambda_{1}$ is either at most $s_{1}$ or at least $s_{0}$ and we further show that $s_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$ and $s_{0}\geq\alpha n-\frac{1+O(\alpha)}{2\alpha^{2}}$ when $n$ is sufficiently large. Therefore, just like in the proof of Theorem 1.2, we have two cases: either $\lambda_{1}$ is small (at most $\frac{1-\alpha^{2}}{\alpha^{2}}$) or large (nearly $\alpha n$)
If $\lambda_{1}$ is small, i.e. $\lambda_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$, then using a straightforward generalization of Lemma 3.2, we will conclude that $n\leq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$. Otherwise, $\lambda_{1}$ is large, i.e. $\lambda_{1}\geq\alpha n-\frac{1+O(\alpha)}{2\alpha^{2}}$, in which case we will proceed as in the proof of Theorem 1.2 by applying Lemma 3.5 with $L=M-\lambda_{1}xx^{*}$ where $x$ is a unit eigenvector corresponding to $\lambda_{1}$ in order to conclude that
$n\leq\frac{1+O(\alpha)}{\alpha}r.\qed$
We first establish the required generalization of Lemma 3.2.
Lemma 4.1.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $S^{1}(\alpha)$-code. If the largest eigenvalue $\lambda_{1}$ of the Gram matrix $M=M_{\mathscr{C}}$ satisfies $\lambda_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$, then
$|\mathscr{C}|\leq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$
with equality only if the span of $\mathscr{C}$ has dimension $11/\alpha^{2}-$ 1.
Proof.
As in the proof of Lemma 3.2, we let $n=|\mathscr{C}|$ and $r=\operatorname{rk}(M)$ and observe that the desired bound follows from
$n(\alpha^{2}n+1-\alpha^{2})=\operatorname{tr}(M^{2})=\sum_{i=1}^{r}{\lambda_{i}(M)^{2}}\leq\lambda_{1}\operatorname{tr}(M)\leq\frac{1-\alpha^{2}}{\alpha^{2}}n.$
Moreover, equality occurs only if $\lambda_{i}(M)=\frac{1-\alpha^{2}}{\alpha^{2}}$ for all $1\leq i\leq r$. By changing bases, we may assume that $\mathscr{C}\subset\mathbb{C}^{r}$, so that if we let $V\in\mathbb{C}^{r\times\mathscr{C}}$ be the matrix given by $Vy=\sum_{v\in\mathscr{C}}{y(v)v}$, then $M=V^{*}V$ has the same nonzero eigenvalues as $VV^{*}=\sum_{v\in\mathscr{C}}{vv^{*}}$. Therefore, $n=\left(1/\alpha^{2}-1\right)^{2}$ implies that $\sum_{v\in\mathscr{C}}{vv^{*}}=\frac{1-\alpha^{2}}{\alpha^{2}}I$ where $I$ is the $r\times r$ identity matrix and by taking the trace, we conclude that $\left(1/\alpha^{2}-1\right)^{2}=n=\frac{1-\alpha^{2}}{\alpha^{2}}r$ and so $r=1/\alpha^{2}-1$. ∎
We now obtain the aforementioned degree 3 polynomial inequality as a function of the largest eigenvalue of the corresponding Gram matrix.
Lemma 4.2.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $S^{1}(\alpha)$-code in $\mathbb{C}^{r}$ with $n=|\mathscr{C}|$ whose Gram matrix $M=M_{\mathscr{C}}$ has largest eigenvalue $\lambda_{1}=\lambda_{1}(M)$. Then we have
$\lambda_{1}^{3}-\alpha^{2}\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)^{2}\lambda_{1}+(1-\alpha^{2})\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)n\geq 0,$
with equality if $n=r^{2}$.
Proof.
Let $x$ be a unit eigenvector corresponding to $\lambda_{1}$. For each $u\in\mathscr{C}$, Theorem 2.4 yields
$(1-\alpha^{2})\lambda_{1}+\frac{\alpha^{2}\lambda_{1}^{4}|x(u)|^{2}}{\alpha^{2}n+1-\alpha^{2}}\geq\alpha^{2}\lambda_{1}^{2}+(1-\alpha^{2})\lambda_{1}^{2}|x(u)|^{2}.$
Since $\alpha,\lambda_{1}>0$ and $x$ is a unit vector, summing the above over all $u\in\mathscr{C}$ and multiplying by $\frac{\alpha^{2}n+1-\alpha^{2}}{\alpha^{2}\lambda_{1}}$ yields the desired inequality. Moreover, if $n=r^{2}$ then via Theorem 2.4 we have equality. ∎
In order to bound the largest root of the degree 3 polynomial in the preceding lemma, we need the following computational lemma.
Lemma 4.3.
Let $p(x)=x^{3}-bx+c$ be a polynomial such that $b>0$, $c\geq 0$, and $\frac{c}{b^{3/2}}\leq\frac{2}{3\sqrt{3}}$. Then the largest root of $p$ is at least $\sqrt{b}-\frac{c}{2b}-\left(\frac{27}{4}-3\sqrt{3}\right)\frac{c^{2}}{b^{5/2}}$.
Proof.
Let $q=\frac{27}{4}-3\sqrt{3}$, $z=\frac{c}{b^{3/2}}$, and $x=1-\frac{z}{2}-qz^{2}$ so that $\sqrt{b}x=\sqrt{b}-\frac{c}{2b}-q\frac{c^{2}}{b^{5/2}}$. Since the leading term of $p$ is positive, we need to show that $p\left(\sqrt{b}x\right)\leq 0$. To this end, we observe that the polynomial $g(z)=\frac{3}{4}-2q+\left(3q-\frac{1}{8}\right)z+\left(3q^{2}-\frac{3q}{4}\right)z^{2}-\frac{3q^{2}}{2}z^{3}-q^{3}z^{4}$ satisfies
$\frac{p\left(\sqrt{b}x\right)}{b^{3/2}}=x^{3}-x+z=z^{2}g(z),$
so we equivalently need to show that $g(z)\leq 0$ for all $0\leq z\leq\frac{2}{3\sqrt{3}}$. To verify this fact, we compute [^12] Although these computations can all be done by hand, we recommend that the reader make use of a mathematical software such as Matlab, Mathematica, or WolframAlpha. that $g\left(\frac{2}{3\sqrt{3}}\right)=0$, so it suffices to check that $g^{\prime}(z)>0$ for all $0\leq z\leq\frac{2}{3\sqrt{3}}$. To see this, we compute that $g^{\prime}(0)>0$, $g^{\prime}(-1)<0$, $g^{\prime}\left(\frac{2}{3\sqrt{3}}\right)>0$ and note that $g^{\prime}$ is a degree 3 polynomial with a negative leading term, so that it must have roots $r_{0},r_{1},r_{2}$ satisfying $r_{2}<-1<r_{1}<0$ and $r_{0}>\frac{2}{3\sqrt{3}}$. ∎
Remark 4.4.
Note that the discriminant of $x^{3}-bx+c$ is $\frac{c^{2}}{4}-\frac{b^{3}}{27}$, which is at most 0 precisely when $\frac{c}{b^{3/2}}\leq\frac{2}{3\sqrt{3}}$, in which case Viète’s formula implies that the largest root of this polynomial will be $2\sqrt{\frac{b}{3}}\cos\left(\frac{1}{3}\arccos\left(\frac{-3\sqrt{3}c}{2b^{3/2}}\right)\right)$. However, to make use of this expression, we would still need to approximate it by a Taylor polynomial with an explicit bound on the error term, so we found it simpler to derive such an approximation directly. Also, the coefficient $\frac{27}{4}-3\sqrt{3}$ is best possible.
We now obtain the desired bounds on the roots of the polynomial from Lemma 4.2. In particular, it will be convenient to obtain a bound on the square of the largest root, as follows.
Lemma 4.5.
Let $10<\alpha<$ 1 and $n\in\mathbb{N}$ be such that $n\geq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$. Then the polynomial $p(x)=x^{3}-\alpha^{2}\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)^{2}x+(1-\alpha^{2})\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)n$ has 3 real roots $s_{0}\geq s_{1}\geq s_{2}$ satisfying $s_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$ and $s_{0}^{2}>n(\alpha^{2}n+1-\alpha^{2})-\frac{(1-\alpha^{2})(1-\alpha)}{\alpha}n-3\frac{(1-\alpha^{2})^{2}}{\alpha^{4}}$.
Proof.
Let $b=\alpha^{2}\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)^{2}$ and $c=(1-\alpha^{2})\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)n$, so that $p(x)=x^{3}-bx+c$. We would like to apply Lemma 4.3, so we need to bound $\frac{c}{b^{3/2}}=\frac{(1-\alpha^{2})n}{\alpha^{3}\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)^{2}}$. One can verify by differentiating this expression with respect to $n$, that it is decreasing when $n\geq\frac{1-\alpha^{2}}{\alpha^{2}}$ and we have $n\geq\max\left(\left(\frac{1}{\alpha^{2}}-1\right)^{2},1\right)\geq\frac{1-\alpha^{2}}{\alpha^{2}}$, so $\frac{c}{b^{3/2}}$ is maximized when $n=\max\left(\left(\frac{1}{\alpha^{2}}-1\right)^{2},1\right)$. Furthermore, observe that for both $n=\left(\frac{1}{\alpha^{2}}-1\right)^{2}$ and $n=1$, we have $\frac{c}{b^{3/2}}=\alpha(1-\alpha^{2})$, so that we may conclude
$\frac{c}{b^{3/2}}\leq\alpha(1-\alpha^{2})\leq\frac{2}{3\sqrt{3}},$
where the last inequality holds since $\alpha(1-\alpha^{2})$ is maximized when $\alpha=1/\sqrt{3}$. We may now apply Lemma 4.3 to conclude that $s_{0}\geq\sqrt{b}-\frac{c}{2b}-\left(\frac{27}{4}-3\sqrt{3}\right)\frac{c^{2}}{b^{5/2}}$ and since $32\left(\frac{27}{4}-3\sqrt{3}\right)-1/4<$ 3, we obtain the desired bound on the largest root
$s_{0}^{2}>b-\frac{c}{\sqrt{b}}-3\frac{c^{2}}{b^{2}}$ $=n(\alpha^{2}n+1-\alpha^{2})+(1-\alpha^{2})n+\frac{\left(1-\alpha^{2}\right)^{2}}{\alpha^{2}}-\frac{1-\alpha^{2}}{\alpha}n-3\frac{(1-\alpha^{2})^{2}n^{2}}{\alpha^{4}\left(n+\frac{1-\alpha^{2}}{2\alpha^{2}}\right)^{2}}$
$>n(\alpha^{2}n+1-\alpha^{2})-\frac{(1-\alpha^{2})(1-\alpha)}{\alpha}n-3\frac{(1-\alpha^{2})^{2}}{\alpha^{4}}.$
Now we compute that $p(0)>0$ and since $n\geq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$, we also compute that $p\left(\frac{1-\alpha^{2}}{\alpha^{2}}\right)\leq 0$. Therefore, as $p$ has a positive leading term, its smallest root must satisfy $s_{2}<0$ and its largest root must satisfy $s_{0}\geq\frac{1-\alpha^{2}}{\alpha^{2}}$. Hence, either $s_{0}=\frac{1-\alpha^{2}}{\alpha^{2}}$ or else $p$ has a root in the interval $\left[0,\frac{1-\alpha^{2}}{\alpha^{2}}\right]$ which is not $s_{0}$ or $s_{2}$, so it must be $s_{1}$. In either case, we conclude the other desired bound $s_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$. ∎
Proof of Theorem 1.11.
We let $\mathscr{C}$ be a spherical $S^{1}(\alpha)$-code representing a set of $n=N^{\mathbb{C}}_{\alpha}(r)$ equiangular lines in $\mathbb{C}^{r}$ and let $M=M_{\mathscr{C}}$ be the corresponding Gram matrix with largest eigenvalue $\lambda_{1}=\lambda_{1}(M)$. Lemma 4.2 implies that $p(\lambda_{1})=\lambda_{1}^{3}-\alpha^{2}\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)^{2}\lambda_{1}+(1-\alpha^{2})\left(n+\frac{1-\alpha^{2}}{\alpha^{2}}\right)n\geq 0$. Just like the proof of Theorem 1.2, we will show that exactly one of the following holds: either $n\leq(1/\alpha^{2}-1)^{2}$ with equality only if $\mathscr{C}$ forms a SIC in $\mathbb{C}^{1/\alpha^{2}-1}$ or $n\leq\frac{1+\alpha}{\alpha}r+3\frac{1+\alpha}{\alpha^{3}}$. Thus we may assume without loss of generality that $n\geq\left(\frac{1}{\alpha^{2}}-1\right)^{2}$ and therefore apply Lemma 4.5 to conclude that $p$ has real roots $s_{0}>s_{1}>s_{2}$ satisfying $s_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$ and $s_{0}^{2}>n(\alpha^{2}n+1-\alpha^{2})-\frac{(1-\alpha^{2})(1-\alpha)}{\alpha}n-3\frac{(1-\alpha^{2})^{2}}{\alpha^{4}}$. Since $p$ has a positive leading term, $p(\lambda_{1})\geq 0$ implies that either $\lambda_{1}\leq s_{1}$ or $\lambda_{1}\geq s_{0}$. If $\lambda_{1}\leq s_{1}\leq\frac{1-\alpha^{2}}{\alpha^{2}}$, then the first desired bound $n\leq\left(1/\alpha^{2}-1\right)^{2}$ and a characterization of equality follow from Lemma 4.1.
Otherwise, we have $\lambda_{1}\geq s_{0}$. Now let $x$ be a unit eigenvector corresponding to $\lambda_{1}$ and let $L=M-\lambda_{1}xx^{*}$. Lemma 3.4 implies that $\operatorname{tr}(L)=n-\lambda_{1}\geq(1-\alpha)(n-1)$ and we also have $\operatorname{rk}(L)=r-1$ and $\operatorname{tr}(L^{2})=\operatorname{tr}(M^{2})-\lambda_{1}^{2}\leq n(\alpha^{2}n+1-\alpha^{2})-s_{0}^{2}$. Therefore, Lemma 3.5 yields
$(1-\alpha)^{2}(n-1)^{2}\leq\operatorname{tr}(L)^{2}\leq(r-1)\operatorname{tr}(L^{2})$ $\leq(r-1)(n(\alpha^{2}n+1-\alpha^{2})-s_{0}^{2})$
$<(r-1)\left(\frac{(1-\alpha^{2})(1-\alpha)}{\alpha}n+3\frac{(1-\alpha^{2})^{2}}{\alpha^{4}}\right)$
and dividing by $(1-\alpha)^{2}n$, we conclude
$n-2<\frac{(n-1)^{2}}{n}<\frac{1+\alpha}{\alpha}(r-1)+3\frac{(1+\alpha)^{2}}{\alpha^{4}}\frac{r}{n}.$
The second desired bound $n\leq\frac{1+\alpha}{\alpha}r+3\frac{1+\alpha}{\alpha^{3}}$ now follows from the above inequality. Indeed, we may assume that $n\geq\frac{1+\alpha}{\alpha}r$ as otherwise we are done, and so we have
$n<\frac{1+\alpha}{\alpha}r-\frac{1+\alpha}{\alpha}+2+3\frac{(1+\alpha)^{2}}{\alpha^{4}}\frac{r}{n}\leq\frac{1+\alpha}{\alpha}r+3\frac{1+\alpha}{\alpha^{3}}.\qed$
5 Eigenvalues of regular graphs
In this subsection, we will prove Theorem 1.15 and Corollary 1.17, which can be seen as generalizations of the Alon–Boppana theorem to dense graphs. Our approach will be to use (2) in order to convert the given regular graph into a corresponding system of real equiangular lines and then apply the projection method of Section 2. As a consequence, our new bounds will be tight for any strongly regular graph corresponding to $\binom{r+1}{2}$ equiangular lines in $\mathbb{R}^{r}$. To this end, we first recall some well-known facts about the second eigenvalue of connected graphs, see e.g. [26] p. 163.
Lemma 5.1.
$\lambda_{2}(K_{n})=-1$ for any complete graph $K_{n}$, $\lambda_{2}(K_{n_{1},\ldots,n_{r}})=0$ for any complete multipartite graph $K_{n_{1},\ldots,n_{r}}$, and $\lambda_{2}(H)>0$ for any other nontrivial connected graph $H$.
Remark 5.2.
Beyond Lemma 5.1, there is a line of research characterizing graphs $H$ with $\lambda_{2}(H)\leq c$ or $\lambda_{2}(H)=c$ for small constant $c$, see [25]. There are also important results on characterizing graphs $H$ with the smallest eigenvalue $\lambda_{n}(H)\geq-c$ for small $c$, such as the classical result for $c=2$ [18, 45] and more recent generalizations [64, 60].
We now derive an improved version of Theorem 2.1 in the special case where $y=\mathbbm{1}$ is an eigenvector of $M$.
Theorem 5.3.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with corresponding Gram matrix $M=M_{\mathscr{C}}$. If $n=|\mathscr{C}|$ and $\mathbbm{1}$ is an eigenvector of $M$ with corresponding eigenvalue $\lambda\neq 0$, then
$\frac{(1-\alpha^{2})n}{2\lambda}\left\langle x,Mx\right\rangle+\left(\frac{\lambda^{2}}{n}-\frac{1-\alpha^{2}}{2}\right)\left\langle x,\mathbbm{1}\right\rangle^{2}\geq\left\langle x,M^{2}x\right\rangle$
for all $x\in\mathbb{R}^{\mathscr{C}}$, with equality whenever $n=\binom{r+1}{2}-1$.
Proof.
Let us suppose we are in the setting of the proof of Theorem 2.1, so that $\mathscr{C}=\{v_{1},\ldots,v_{n}\}$, the linear map $\mathscr{W}\colon\mathbb{R}^{\mathscr{C}}\rightarrow\mathscr{S}_{r}$ is defined by $\mathscr{W}e_{v}=vv^{\intercal}$ for $v\in\mathscr{C}$ and the orthogonal projection $\mathscr{P}\colon\mathscr{S}_{r}\rightarrow\mathscr{S}_{r}$ onto the span of $\{vv^{\intercal}:v\in\mathscr{C}\}$ is given by $\mathscr{P}=\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}$.
Now let $x\in\mathbb{R}^{\mathscr{C}}$ and define $X_{x}=\frac{1}{2}\left(Vx(V\mathbbm{1})^{\intercal}+V\mathbbm{1}(Vx)^{\intercal}\right)$. We compute that
$||X_{x}||_{F}^{2}=\frac{1}{2}\left(\left\langle x,Mx\right\rangle\left\langle\mathbbm{1},M\mathbbm{1}\right\rangle+\left\langle x,M\mathbbm{1}\right\rangle^{2}\right)=\frac{1}{2}\left(\lambda n\left\langle x,Mx\right\rangle+\lambda^{2}\left\langle x,\mathbbm{1}\right\rangle^{2}\right).$ (4)
Moreover, for all $v\in\mathscr{C}$, we have $\left\langle vv^{\intercal},V\mathbbm{1}(Vx)^{\intercal}\right\rangle_{F}=\left\langle e_{v},M\mathbbm{1}\right\rangle\left\langle e_{v},Mx\right\rangle=\lambda(Mx)(v)$ and similarly $\left\langle vv^{\intercal},V\mathbbm{1}(Vx)^{\intercal}\right\rangle_{F}=\lambda(Mx)(v)$, so that $\mathscr{W}^{\#}X_{x}=\lambda Mx$. As in the proof of Theorem 2.1, we have $(\mathscr{W}^{\#}\mathscr{W})^{-1}=\frac{1}{1-\alpha^{2}}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)$ and so
$||\mathscr{P}X_{x}||_{F}^{2}=X_{x}^{\#}\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}X_{x}$ $=\frac{\lambda^{2}}{1-\alpha^{2}}(Mx)^{\intercal}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)Mx$ (5)
$=\frac{\lambda^{2}}{1-\alpha^{2}}\left(\left|\left|Mx\right|\right|^{2}-\frac{\alpha^{2}\lambda^{2}}{\alpha^{2}n+1-\alpha^{2}}\left\langle x,\mathbbm{1}\right\rangle^{2}\right).$
In particular, applying (4) and (5) with $x=\mathbbm{1}$ yields $||X_{\mathbbm{1}}||_{F}^{2}=\lambda^{2}n^{2}$ and
$||\mathscr{P}X_{\mathbbm{1}}||_{F}^{2}=\frac{1}{1-\alpha^{2}}\left(\lambda^{4}n-\frac{\alpha^{2}\lambda^{4}n^{2}}{\alpha^{2}n+1-\alpha^{2}}\right)=\frac{\lambda^{4}n}{\alpha^{2}n+1-\alpha^{2}}.$
We now claim that $X_{\mathbbm{1}}-\mathscr{P}X_{\mathbbm{1}}\neq 0$. Indeed, otherwise we would have
$0=||X_{\mathbbm{1}}-\mathscr{P}X_{\mathbbm{1}}||_{F}^{2}=||X_{\mathbbm{1}}||_{F}^{2}-||\mathscr{P}X_{\mathbbm{1}}||_{F}^{2}=\lambda^{2}n\left(n-\frac{\lambda^{2}}{\alpha^{2}n+1-\alpha^{2}}\right),$
which implies $\lambda^{2}=n(\alpha^{2}n+1-\alpha^{2})$. However, we also have $\sum_{i=1}^{n}{\lambda_{i}(M)^{2}}=\operatorname{tr}(M^{2})=n(\alpha^{2}n+1-\alpha^{2})$, so that all eigenvalues of $M$ except $\lambda$ must be 00 and thus $M=\frac{\lambda}{n}\mathbbm{1}\mathbbm{1}^{\intercal}$, which contradicts the fact that $M(v,v)=1$ and $|M(u,v)|=\alpha<1$ for any $u\neq v\in\mathscr{C}$.
Therefore, we may define
$Z=\frac{X_{\mathbbm{1}}-\mathscr{P}X_{\mathbbm{1}}}{\sqrt{||X_{\mathbbm{1}}||_{F}^{2}-||\mathscr{P}X_{\mathbbm{1}}||_{F}^{2}}}=\frac{X_{\mathbbm{1}}-\mathscr{P}X_{\mathbbm{1}}}{\lambda\sqrt{n^{2}-\frac{\lambda^{2}n}{\alpha^{2}n+1-\alpha^{2}}}}$
so that by definition, $Z$ is orthogonal to $vv^{\intercal}$ for all $v\in\mathscr{C}$. It follows that $\mathscr{P}^{\prime}=\mathscr{P}+ZZ^{\#}$ is the orthogonal projection onto the span of $\{vv^{\intercal}:v\in\mathscr{C}\}\cup Z$ and moreover,
$||X_{x}||_{F}^{2}\geq||\mathscr{P}^{\prime}X_{x}||_{F}^{2}=||\mathscr{P}X_{x}||_{F}^{2}+\left\langle Z,X_{x}\right\rangle_{F}^{2}.$ (6)
In view of (4) and (5), it remains to compute $\left\langle Z,X_{x}\right\rangle_{F}$. To this end, observe that
$\left\langle X_{\mathbbm{1}},Vx(V\mathbbm{1})^{\intercal}\right\rangle_{F}=\left\langle V\mathbbm{1},Vx\right\rangle\left\langle V\mathbbm{1},V\mathbbm{1}\right\rangle=\left\langle M\mathbbm{1},x\right\rangle\left\langle\mathbbm{1},M\mathbbm{1}\right\rangle=\lambda^{2}n\left\langle x,\mathbbm{1}\right\rangle$
and similarly $\left\langle X_{\mathbbm{1}},V\mathbbm{1}(Vx)^{\intercal}\right\rangle_{F}=\lambda^{2}n\left\langle x,\mathbbm{1}\right\rangle$, so that $\left\langle X_{\mathbbm{1}},X_{x}\right\rangle_{F}=\lambda^{2}n\left\langle x,\mathbbm{1}\right\rangle$. Since $\mathscr{W}^{\#}X_{\mathbbm{1}}=\lambda M\mathbbm{1}=\lambda^{2}\mathbbm{1}$, we conclude that
$\left\langle\mathscr{P}X_{\mathbbm{1}},X_{x}\right\rangle=X_{\mathbbm{1}}^{\#}\mathscr{W}(\mathscr{W}^{\#}\mathscr{W})^{-1}\mathscr{W}^{\#}X_{x}$ $=\frac{\lambda^{3}}{1-\alpha^{2}}\mathbbm{1}^{\intercal}\left(I-\frac{\alpha^{2}}{\alpha^{2}n+1-\alpha^{2}}J\right)Mx$
$=\frac{\lambda^{3}}{1-\alpha^{2}}\left(1-\frac{\alpha^{2}n}{\alpha^{2}n+1-\alpha^{2}}\right)\mathbbm{1}^{\intercal}Mx$
$=\frac{\lambda^{4}}{\alpha^{2}n+1-\alpha^{2}}\left\langle x,\mathbbm{1}\right\rangle$
and so we have
$\left\langle Z,X_{x}\right\rangle_{F}=\frac{\left\langle X_{\mathbbm{1}},X_{x}\right\rangle_{F}-\left\langle\mathscr{P}X_{\mathbbm{1}},X_{x}\right\rangle_{F}}{\lambda\sqrt{n^{2}-\frac{\lambda^{2}n}{\alpha^{2}n+1-\alpha^{2}}}}$ $=\frac{\lambda^{2}n\left\langle x,\mathbbm{1}\right\rangle-\frac{\lambda^{4}}{\alpha^{2}n+1-\alpha^{2}}\left\langle x,\mathbbm{1}\right\rangle}{\lambda\sqrt{n^{2}-\frac{\lambda^{2}n}{\alpha^{2}n+1-\alpha^{2}}}}$
$=\frac{\lambda}{n}\sqrt{n^{2}-\frac{\lambda^{2}n}{\alpha^{2}n+1-\alpha^{2}}}\left\langle x,\mathbbm{1}\right\rangle.$
Now that we have determined $\left\langle Z,X_{x}\right\rangle_{F}$, we conclude via (4), (5), and (6) that
$\frac{\lambda n}{2}\left\langle x,Mx\right\rangle\geq\frac{\lambda^{2}}{1-\alpha^{2}}\left\langle x,M^{2}x\right\rangle-\left(\frac{\lambda^{4}}{(1-\alpha^{2})n}-\frac{\lambda^{2}}{2}\right)\left\langle x,\mathbbm{1}\right\rangle^{2},$
which is equivalent to the desired bound.
Moreover, note that $\mathscr{S}_{r}$ has dimension $\binom{r+1}{2}$ so that if $n=\binom{r+1}{2}-1$, then since $\operatorname{rk}(\mathscr{W})=n$ and $Z$ is orthogonal to $vv^{\intercal}$ for all $v\in\mathscr{C}$, we have that $\{vv^{\intercal}:v\in\mathscr{C}\}\cup Z$ span $\mathscr{S}_{r}$ and thus $\mathscr{P}^{\prime}$ is the identity map, giving equality above. ∎
Remark 5.4.
The inequality of Theorem 5.3 is equivalent to the matrix $\frac{1-\alpha^{2}}{2\lambda}M+\left(\frac{\lambda^{2}}{n}-\frac{1-\alpha^{2}}{2}\right)J-M^{2}$ being positive semidefinite.
Lemma 5.5.
Let $10<\alpha<$ 1 and let $\mathscr{C}$ be a spherical $\{\alpha,-\alpha\}$-code in $\mathbb{R}^{r}$ with corresponding Gram matrix $M=M_{\mathscr{C}}$. If $n=|\mathscr{C}|$ and $\mathbbm{1}$ is an eigenvector of $M$ with corresponding eigenvalue $\lambda\neq 0$, then
$\lambda^{2}\geq n(\alpha^{2}n+1-\alpha^{2})-\frac{1-\alpha^{2}}{2}n\left(\frac{n}{\lambda}-1\right),$
and for any eigenvalue $\mu$ of $M$ with an eigenvector orthogonal to $\mathbbm{1}$, we have
$\mu\leq\frac{(1-\alpha^{2})n}{2\lambda},$
with equality in both whenever $n=\binom{r+1}{2}-1$.
Proof.
Fix any $v\in\mathscr{C}$ and observe that $\left\langle e_{v},M^{2}e_{v}\right\rangle=\alpha^{2}n+1-\alpha^{2}$. Thus applying Theorem 5.3 with $x=e_{v}$, we obtain
$\frac{1-\alpha^{2}}{2}\left(\frac{n}{\lambda}-1\right)+\frac{\lambda^{2}}{n}=\frac{(1-\alpha^{2})n}{2\lambda}+\frac{\lambda^{2}}{n}-\frac{1-\alpha^{2}}{2}\geq\alpha^{2}n+1-\alpha^{2},$
which is equivalent to the first bound.
For the second bound, if $\mu=0$ then it holds trivially. Otherwise, we let $x$ be a unit eigenvector for $\mu$. Since $\left\langle x,\mathbbm{1}\right\rangle=0$, we have via Theorem 5.3 that $\frac{(1-\alpha^{2})n}{2\lambda}\mu\geq\mu^{2}$, so dividing by $\mu$ gives the desired bound. Moreover, $n=\binom{r+1}{2}-1$ implies we have equality in Theorem 5.3 and therefore also in both of the above bounds. ∎
Using Lemma 5.5, we are now able to obtain the desired bounds on the eigenvalues of regular graphs and prove Theorem 1.15.
Proof of Theorem 1.15.
First note that if $k=0$ then both inequalities hold trivially. Assume now that $k\geq 1$. We claim that $\lambda_{2}>0$ follows from Lemma 5.1. Indeed, otherwise $G$ would have to either be a complete graph, a complete multipartite graph, or disconnected. Moreover, $G$ cannot be the complete graph or a regular complete multipartite graph, since they all have spectral gap at least $n/2$. Otherwise, $G$ isn’t connected, in which case the Perron–Frobenius theorem implies $\lambda_{2}=k\geq 1$.
Now let $A=A(G)$ be the adjacency matrix of $G$ and for each $i\in[n]$, let $\lambda_{i}=\lambda_{i}(G)$ be the $i$th largest eigenvalue of $G$. By the spectral theorem for symmetric matrices, we have $A=\sum_{i=1}^{n}{\lambda_{i}u_{i}u_{i}^{\intercal}}$ for an orthonormal basis of eigenvectors $u_{1},\ldots,u_{n}\in\mathbb{R}^{n}$. As previously mentioned, the Perron–Frobenius theorem implies $\lambda_{1}=k$ and $u_{1}$ can be taken to be a scalar multiple of $\mathbbm{1}$. Since $||\mathbbm{1}||=\sqrt{n}$ we may take $u_{1}=\frac{1}{\sqrt{n}}\mathbbm{1}$. Since $J=\mathbbm{1}\mathbbm{1}^{\intercal}$, we therefore have
$A=\frac{k}{n}J+\sum_{i=2}^{n}{\lambda_{i}u_{i}u_{i}^{\intercal}}.$
Let $U=(u_{1},\ldots,u_{n})$ be the $n\times n$ matrix with $u_{1},\ldots,u_{n}$ as columns. Since $u_{1},\ldots,u_{n}$ form an orthonormal basis, $U$ is an orthogonal matrix and so we have
$I=UU^{\intercal}=\sum_{i=1}^{n}{u_{i}u_{i}^{\intercal}}=\frac{1}{n}J+\sum_{i=2}^{n}{u_{i}u_{i}^{\intercal}}.$
We now define $\alpha=\frac{1}{2\lambda_{2}+1}$ and note that $10<\alpha<$ 1 since $\lambda_{2}>0$. In view of (2), we define $M=\alpha J-2\alpha A+(1-\alpha)I$ and observe that $M(v,v)=1$ and $|M(u,v)|=\alpha$ for all $u\neq v\in\mathscr{C}$. Moreover, we have that
$M=\frac{1}{2\lambda_{2}+1}\left(J-2A+2\lambda_{2}I\right)=\frac{1}{2\lambda_{2}+1}\left(\left(1-\frac{2(k-\lambda_{2})}{n}\right)J+2\sum_{i=2}^{n}{(\lambda_{2}-\lambda_{i})}u_{i}u_{i}^{\intercal}\right),$
which implies that $M$ is positive semidefinite and has $\mathbbm{1}$ as an eigenvector with corresponding eigenvalue $\lambda=\frac{n-2(k-\lambda_{2})}{2\lambda_{2}+1}>0$. It follows that the dimension of the null space of $M$ is precisely the multiplicity of $\lambda_{2}$ in $G$, $m_{G}(\lambda_{2})$, so that $r=\operatorname{rk}(M)=n-m_{G}(\lambda_{2})$. Therefore, we can associate a vector in $\mathbb{R}^{r}$ to each vertex of $G$, so that the resulting collection $\mathscr{C}$ has Gram matrix $M_{\mathscr{C}}=M$. It follows that $\mathscr{C}$ is a spherical $\{\alpha,-\alpha\}$-code and so we may apply Lemma 5.5 to obtain
$\lambda^{2}$ $\geq n(\alpha^{2}n+1-\alpha^{2})-\frac{(1-\alpha^{2})n^{2}}{2\lambda}+\frac{(1-\alpha^{2})n}{2}$
$=\frac{n(n+(2\lambda_{2}+1)^{2}-1)}{(2\lambda_{2}+1)^{2}}-\frac{((2\lambda_{2}+1)^{2}-1)n^{2}}{2(2\lambda_{2}+1)(n-2(k-\lambda_{2}))}+\frac{(2\lambda_{2}+1)^{2}-1)n}{2(2\lambda_{2}+1)^{2}}.$
Multiplying by $\frac{(2\lambda_{2}+1)^{2}}{n}$, we therefore have
$n-4k+4\lambda_{2}+\frac{4(k-\lambda_{2})^{2}}{n}=\frac{(2\lambda_{2}+1)^{2}}{n}\lambda^{2}\geq n-\frac{2\lambda_{2}(2\lambda_{2}+1)(\lambda_{2}+1)}{1-\frac{2(k-\lambda_{2})}{n}}+6\lambda_{2}^{2}+6\lambda_{2},$
which implies the first desired bound
$2k-\frac{2(k-\lambda_{2})^{2}}{n}\leq\frac{\lambda_{2}(2\lambda_{2}+1)(\lambda_{2}+1)}{1-\frac{2(k-\lambda_{2})}{n}}-\lambda_{2}(3\lambda_{2}+1).$
Moreover, it follows from the spectral decomposition of $M$ that it also has $u_{n}$ as an eigenvector with corresponding eigenvalue $\mu=\frac{2(\lambda_{2}-\lambda_{n})}{2\lambda_{2}+1}$ and so we may apply Lemma 5.5 to obtain
$\frac{2(\lambda_{2}-\lambda_{n})}{2\lambda_{2}+1}=\mu\leq\frac{(1-\alpha^{2})n}{2\lambda}=\frac{((2\lambda_{2}+1)^{2}-1)n}{2(2\lambda_{2}+1)(n-2(k-\lambda_{2}))}.$
Multiplying by $\frac{(2\lambda_{2}+1)}{2}$ and subtracting $\lambda_{2}$ we conclude the second desired bound
$-\lambda_{n}\leq\frac{\lambda_{2}(\lambda_{2}+1)}{1-\frac{2(k-\lambda_{2})}{n}}-\lambda_{2}.$
Finally, $n=\binom{n-m(\lambda_{2})+1}{2}-1=\binom{r+1}{2}-1$ implies we have equality in both bounds of Lemma 5.5 and therefore also in both of the above. ∎
Proof of Corollary 1.17.
We first suppose that there exists a constant $\varepsilon>0$ such that $k-\lambda_{2}\leq(1-\varepsilon)\frac{n}{2}$, or equivalently $1-\frac{2(k-\lambda_{2})}{n}\geq\varepsilon$. Via Theorem 1.15 and Remark 1.16 we have
$k<2\left(k-\frac{(k-\lambda_{2})^{2}}{n}\right)\leq\frac{\lambda_{2}(2\lambda_{2}+1)(\lambda_{2}+1)}{\varepsilon}-\lambda_{2}(3\lambda_{2}+1)\leq O(\lambda_{2}^{3}),$
and therefore $\lambda_{2}\geq\Omega\left(k^{1/3}\right)$. Also via Theorem 1.15, we have $-\lambda_{n}\leq\frac{\lambda_{2}(\lambda_{2}+1)}{\varepsilon}-\lambda_{2}\leq O(\lambda_{2}^{2})$, so that $\lambda_{2}\geq\Omega\left(\sqrt{-\lambda_{n}}\right)$. Moreover, if $G$ is bipartite, then $\lambda_{n}=-k$ and so $\lambda_{2}\geq\Omega\left(\sqrt{k}\right)$.
Now let us further suppose that $k-\lambda_{2}=o(n)$. In this case, Theorem 1.15 implies
$(1-o(1))2k+o(1)\lambda_{2}=2\left(k-\frac{(k-\lambda_{2})^{2}}{n}\right)\leq\frac{\lambda_{2}(2\lambda_{2}+1)(\lambda_{2}+1)}{1-o(1)}-\lambda_{2}(3\lambda_{2}+1)$ $=(1+o(1))2\lambda_{2}^{3},$
so that $\lambda_{2}\geq(1-o(1))k^{1/3}$. Again using Theorem 1.15 we have $-\lambda_{n}\leq\frac{\lambda_{2}(\lambda_{2}+1)}{1-o(1)}-\lambda_{2}\leq(1+o(1))\lambda_{2}^{2}$, so that $\lambda_{2}\geq(1-o(1))\sqrt{-\lambda_{n}}$. Finally, if $G$ is bipartite, we have $\lambda_{n}=-k$ and so $\lambda_{2}\geq(1-o(1))\sqrt{k}$. ∎
6 Concluding remarks
In this section, we make some concluding remarks and suggest some directions for future research.
  1. 1.
    Regarding real equiangular lines, one of the most interesting questions is to determine how far the approach of Jiang, Tidor, Yao, Zhang, and Zhao [61] can be extended. Their method relied on two ingredients: A bound on the maximum degree $\Delta$ of a corresponding graph $G$ which only depends on $\alpha$, and a sublinear bound on the multiplicity $m(\lambda_{2})$ of the second eigenvalue $\lambda_{2}$ of any connected graph with maximum degree at most $\Delta$. Since we now have strong bounds on $\Delta$, the only limitation to extending their result is that their bound on second eigenvalue multiplicity $m(\lambda_{2})$ for a connected graph with maximum degree $\Delta$ is only $O\left(n\log{\Delta}/\log{\log{n}}\right)$. Note that for normalized adjacency matrices and therefore also for regular graphs, McKenzie, Rasmussen, and Srivastava [71] improve on the approach of [61] to improve the upper bound to $O_{\Delta}(n/\log^{c}{n})$ for some constant $c$. However, they give reason to suggest that their methods cannot go beyond $O_{\Delta}\left(n/\log{n}\right)$. Moreover, Haiman, Schildkraut, Zhang, and Zhao [52] give a construction of a graph with max degree $\Delta=4$ and $m(\lambda_{2})\geq\sqrt{n/\log{n}}$, and they also point out that for bounded degree ($\Delta=O(1)$) graphs, $\sqrt{n}$ is a natural barrier for group representation based constructions such as theirs.
  2. 2.
    Regarding the multiplicity of the second eigenvalue of a graph, we also point out that a related question was considered by Colin de Verdiére [87]. He conjectured that the maximum of the multiplicity of the second smallest eigenvalue of a generalized Laplacian operator, over all such operators on an orientable surfaces $S$, is precisely $\operatorname{chr}(S)-1$ where $\operatorname{chr}(S)$ is the chromatic number of the surface. Moreover, via Heawood’s formula [78] it is known that $\operatorname{chr}(S)=\left\lfloor\left(7+\sqrt{48\text{g}(S)+1}\right)/2\right\rfloor$ where $\text{g}(S)$ is the genus of $S$. In view of this, Tlusty [86] formulates an analogous statement in the setting of graphs: Given a graph $G$, the maximum multiplicity of the second eigenvalue of a weighted [^13] For a graph $G$ with vertex set $V(G)=[n]$, a symmetric matrix $L\in\mathbb{R}^{n\times n}$ is called a weighted Laplacian of $G$ if $L\mathbbm{1}=0$ and for all $i\neq j$, $L_{i,j}=0$ when $ij\notin E(G)$ and $L_{i,j}<0$ when $ij\in E(G)$. Laplacian matrix on $G$ is $\left\lfloor\left(7+\sqrt{48\text{g}(G)+1}\right)/2\right\rfloor-1$, where $g(G)$ is the minimal genus of a surface in which $G$ embeds. Moreover, he proves this claim for a class of graphs including paths, cycles, complete graphs, and their Cartesian products. It is not hard [^14] Embed the vertices $V(G)$ arbitrarily on the sphere and for each edge $uv\in E(G)$, add a handle near $u$ and $v$ and pass the edge along this handle. to see that $g(G)\leq|E(G)|$ and thus if Tlusty’s statement were to be true for say, the unweighted Laplacian $kI-A$ of a $k$-regular graph $G$ on $n$ vertices with adjacency matrix $A$, then we would be able to conclude that $m_{G}(\lambda_{2})\leq O\left(\sqrt{|E(G)|}\right)=O\left(\sqrt{kn}\right)$. Also note that for graphs whose corresponding matrix satisfies a certain technical condition known as the strong Arnold hypothesis, Theorem 5 of Pendavingh [74] implies the asymptotic version of this result $m_{G}(\lambda_{2})\leq O(\sqrt{|E(G)|})$. In particular, for bounded degree graphs this give a bound of $O(\sqrt{n})$, matching the natural barrier discussed previously.
  3. 3.
    From the point of view of applications, one of the most important questions regarding complex equiangular is Zauner’s conjecture, i.e. that there exist $r^{2}$ equiangular lines (a SIC) in $\mathbb{C}^{r}$ for all $r$. An interesting consequence of Theorem 1.11 is that if one constructs $r^{2}$ equiangular lines in $\mathbb{C}^{(1-o(1))(r+1)^{3/2}}$ with $\alpha=1/\sqrt{r+1}$, then one immediately obtains a SIC in $\mathbb{C}^{r}$, thereby providing a potential alternative approach to proving Zauner’s conjecture.
  4. 4.
    Another natural question is whether the graph-based methods used in the real case can be extended to complex equiangular lines. We remark that starting with the Gram matrix $M$ of a spherical $S^{1}(\alpha)$-code $\mathscr{C}$ in $\mathbb{C}^{r}$, one may use (2) to define a matrix $A$ which is like a complex generalization of an adjacency matrix, and moreover that our approach can be generalized to give an upper bound on $\max_{v\in\mathscr{C}}{\text{Re}\left(e_{v}^{\intercal}A{\mathbbm{1}}\right)}$, a quantity analogous to the maximum degree. However, we do not know of an appropriate Alon–Boppana-type theorem in this setting.
  5. 5.
    Since we have a way of generalizing the maximum degree bound to complex equiangular lines, it would be interesting to determine if the results of Jiang, Tidor, Yao, Zhang, and Zhao [61] can be extended in order to exactly determine $N^{\mathbb{C}}_{\alpha}(r)$ when $r$ is large relative to $1/\alpha$. We note that this would require complex versions of the bound on the multiplicity of the second largest eigenvalue, as well as a complex version of the Perron–Frobenius theorem.
  6. 6.
    We observe that the projection method used in this paper can be applied to other matrices associated with a graph, such as the unweighted Laplacian, in order to obtain new spectral inequalities. Indeed, the method applies for any positive semidefinite matrix and any matrix can be made as such by adding a sufficient multiple of the identity.
  7. 7.
    It would be interesting to obtain generalizations of our results to arbitrary spherical $L$-codes in $\mathbb{R}^{r}$ with $|L|=s$, i.e. to $s$-distance sets for $s\geq 2$, as well as for $L=[-1,\alpha]$, i.e. the classical coding theory question which is equivalent to packing spherical caps on a sphere in $\mathbb{R}^{r}$. In particular, as a first step towards understanding $[-1,\alpha]$-codes, it would be interesting to see if our methods can be applied for $[-\alpha,\alpha]$-codes.
  8. 8.
    We expect that our approach should extend beyond lines to higher-order equiangular subspaces with respect to different notions of angle, which are described in [9]. In particular, we predict that our methods can be generalized to equiangular subspaces with respect to the chordal distance.
  9. 9.
    The first Welch bound is the first in a family of higher-order Welch bounds. By generalizing our projection method to higher-order tensors, we expect that it is possible to obtain an improvement to all of the Welch bounds in the same way as we have done for the first.
  10. 10.
    Note that spherical $\{\alpha,-\alpha\}$-codes correspond to signed complete graphs, and so it would be interesting to generalize and apply our methods to signed graphs and more generally, to unitary signings of graphs. In particular, Koolen, Cao, and Yang [63] have recently used a Ramsey-theoretic approach analogous to [7] in order to study signed graphs, and thus we believe that our methods can be applied to extend their result.
Acknowledgments.  We thank Noga Alon and Michael Krivelevich for stimulating discussions, as well as Dustin Mixon for information regarding the novelty of one of our bounds. We also thank Boris Bukh, Yufei Zhao, and Gaston Burrull for helpful feedback on earlier versions of this paper. Finally, we thank Sheila Friedmann for her support during the time this research was being done.